### v1

# Human-Machine Interfaces for Subsea Telerobotics: From Soda-straw to Natural Language Interactions

Abstract
This review explores the evolution of human-machine interfaces (HMIs) for subsea telerobotics, tracing back the transition from traditional first-person “soda-straw” consoles (narrow field-of-view camera feed) to advanced interfaces powered by gesture recognition, virtual reality, and natural language models. First, we discuss various forms of subsea telerobotics applications, current state-of-the-art (SOTA) interface systems, and the challenges they face in robust underwater sensing, real-time estimation, and low-latency communication. Through this analysis, we highlight how advanced HMIs facilitate intuitive interactions between human operators and robots to overcome these challenges. A detailed review then categorizes and evaluates the cutting-edge HMI systems based on their offered features from both human perspectives (e.g., enhancing operator control and situational awareness) and machine perspectives (e.g., improving safety, mission accuracy, and task efficiency). Moreover, we examine the literature on bidirectional interaction and intelligent collaboration in terms of sensory feedback and intuitive control mechanisms for both physical and virtual interfaces. The paper concludes by identifying critical challenges, open research questions, and future directions, emphasizing the need for multidisciplinary collaboration in subsea telerobotics.
keywords:
Subsea telerobotics; marine robotics; human-machine interface; shared autonomy.
1 1. Introduction
[…]
Early teleoperated systems were limited in terms of the feedback they provided, often relying on basic visual data from low-resolution cameras. The operator had to rely heavily on experience and intuition to navigate and perform tasks, which increased the cognitive load and risk of errors. While effective for simpler tasks, studies indicate that the limited field-of-view (FOV), appearing as looking through a “soda straw” [14], results in reduced situational awareness in more complex and dynamic environments, leading to increased cognitive load and reduced performance. The 2D visual feedback provides only a partial understanding of the remote workspace, requiring operators to frequently switch between different viewpoints if available [15], and manually integrate information from various perspectives. This challenge is especially pronounced in confined or cluttered spaces, a common scenario in underwater pipelines, caves, and shipwreck exploration. In larger workspaces, traditional monocular camera-based interfaces fail to provide a comprehensive spatial view, leading to significant cognitive overload on the teleoperators.
These limitations have led to the development of more interactive systems including 3D scene rendering with augmented/virtual/mixed reality (XR), haptic cues, and natural language (NL) controls for complex teleoperation tasks. Recent research highlights the importance of human-robot interface design in improving the operator’s awareness, cognitive load, and stress factors [16, 17, 18] – as well as improving the ROVs’ physical capabilities [19]. These developments have allowed operators to engage in human-machine shared autonomy, significantly enhancing the precision and safety of subsea missions [20, 21, 22]. Transparency and reliability are two other critical attributes of shared autonomy, to provide the operator with clear and interpretable insights into the robot’s intentions and actions [23]. Additionally, fostering mutual trust between the operator and the robot is essential, so that the operator feels both in control and supported by the machine’s autonomous capabilities [24].
Fig. 1 presents a general outline of human-machine sensorimotor interaction for subsea telerobotics. As shown, the interface elements bridge the human and machine entities with an interactive flow of teleoperation (teleop) control and feedback commands. Modern interfaces can integrate virtual engines to render a physics-simulated world and digital twin (DT) models of real ROVs [25, 26]. These virtual scenarios, simulated independently [27] or modeled from real-world data [28], are useful for mission training and rehearsal, as real underwater deployments involve significant costs and logistics [29]. DTs also allow for simulating synthetic sensory data and mission logs, supporting the early-stage development and testing of new perception and navigation algorithms in a controlled environment [30, 31, 32].
Modern subsea ROVs integrate various shared autonomy and semi-autonomous control capabilities, where the robot can handle low-level functions such as station-keeping or obstacle avoidance [33, 22]. This blend of human awareness with machine intelligence reduces cognitive workload and allows the operator to focus on high-level tasks [34], improving both operational safety and efficiency. Customizable interfaces that can adapt to the specific mission needs or the skill level of the operator further streamline the process [35], allowing for greater flexibility and adaptability in a wide range of applications.
Figure 1: An outline of the human-robot sensorimotor interaction flow for subsea telerobotics is shown. The interactions occur via control and feedback devices such as visual console, haptic gloves and suit, HMD, as well as by natural languages. A virtual simulation engine is often integrated with the interface to project the remote environment and ROV operation as a digital twin. Simulators with integrated twin models help teleoperation training by mission rehearsal/replay, which is essential for analyzing teleoperators’ cognitive load as well as for prototyping human-machine embodied/shared autonomy features. (FPV: first person view; NL: natural language; XR: augmented, virtual, and mixed reality)
To this end, contemporary studies emphasize the role of feedback realization, cognitive reasoning, and intuitive control modes for mission safety, efficiency, and success [16, 36]. Multimodal feedback has been particularly useful in engaging teleoperators’ sensations through means such as vibrotactile cues on the palm [37], force and tension on the forearm [38], augmented visual displays [39], and textual responses to queries [40]. On the other hand, control commands are typically propagated by traditional buttons, hand gestures [41], body movements [42], or NL instructions [43]. With any/some of these features, a dynamic feedback-control loop significantly increases the sense of telepresence and enhances mission reliability and success rates of complex subsea tasks.
The interdisciplinary nature of modern interface technologies, encompassing haptics, XR, and NL features – has led to a diverse body of research literature. Fig. 2 illustrates a snapshot of subsea HMI technologies developed across academia and industry. These technologies are often built for specific platforms and applications, revealing an absence of a set of standard guidelines. As a result, the body of knowledge remains scattered, with no comprehensive review that compares the technologies, compiles the findings, and highlights opportunities for integration and standardization.
Figure 2: A chronological evolution of subsea HMI technologies is shown. The top row highlights early achievements and milestones. As technology advanced, this century observed industry solutions for interfaces and simulators, incorporating XR and haptic features. In the last five years, NL-based interactions have gained popularity, driven by the success of LLMs.
In this paper, we compile a comprehensive review of the prominent technologies and relevant literature on subsea HMIs over the last few decades. We also identify the knowledge gaps and highlight current trends to facilitate future research. Specifically, we make the following contributions.
  1. 1.
    We categorize the state-of-the-art (SOTA) subsea telerobotic interfaces and simulators based on attributes such as their choice of sensorimotor interfaces, control and feedback mechanisms, platform compatibility, and other relevant features. We then evaluate and compare these features based on their applicability and effectiveness across diverse subsea applications. These discussions are in Section 2-3.
  2. 2.
    We present a detailed review of human-robot interaction and shared autonomy frameworks in subsea telerobotics, highlighting various design issues regarding bidirectional interaction, degree of intelligence, transparency, trust, and the underlying task allocation processes; see Section 4. These analyses also provide valuable insights into human-machine teaming and the associated feedback engineering and interactive control mechanisms; see Section 5.
  3. 3.
    Finally, we analyze the SOTA approaches and current trends to address the unique challenges in subsea HMIs, highlighting the open problems and potential new use cases for future research; see Section 6.
Through this detailed examination of trends, challenges, and potential breakthroughs – we hope to inspire new approaches to bridge the gap between theory and practice. We anticipate that this review will serve as a comprehensive resource for researchers and engineers and support the development of more efficient and reliable telerobotic systems to advance subsea robotics technologies of the next generation.
2 2  Subsea Telerobotic User Interfaces
Telerobotic interfaces have evolved significantly over the past decades, with novel integrated features playing a critical role in enhancing operator control and mission success rate. Traditionally, vision-based interfaces dominated, where operators relied solely on 2D video feeds to navigate and control a robot [4, 44]. Haptic feedback was integrated later [37], improving task precision by allowing operators to feel forces such as drag from fluid dynamics and resistance from interactions with objects. The introduction of XR interfaces [45] further advanced ROV teleoperation by immersing operators in a 3D virtual environment, providing greater situational awareness and enhancing operator control [46, 47]. More recently, researchers have integrated natural language understanding capability into mission management interfaces [48], allowing operators to plan missions and control ROVs using NL commands, making interaction more intuitive.
Table 1: Comparison of existing subsea teleop interfaces based on their integrated features and add-ons. ($\infty$: unlimited number; W: Windows; M: MacOS; L: Linux).
  Open Video Waypoint Scene 3D Visual Fault Mission Max. # Supported
Source Recorder Navigation Mapper Supported Diagnosis Replay Vehicles OS: W,M,L
  QGroundControl [49] $77$ ✓✓✓
Mission Planner [50] $\infty$ ✓✓✕
APM Planner2 [51] $\infty$ ✓✓✓
EOD Workspace [52] $\infty$ ✓✕✕
VideoRay Cockpit [53] $11$ ✓✕✕
Teledyne SeaBotix [54] $11$ ✓✕✕
SRS Fusion [55] $11$ ✓✕✕
NaviSuite Mobula [56] $11$ ✓✕✕
Abyssal Offshore [57] $11$ ✓✕✕
BRIDGE [58] $11$ ✓✕✕
SubNav [59] $11$ ✓✓✓
DroidPlanner [60] $11$ Android
3 2.1   Status Quo of Robot Teleoperation Interface
Subsea HMIs have been developed to address the unique challenges of underwater operations, where visibility, communication bandwidth, localization techniques, and environmental conditions significantly differ from terrestrial systems. The interfaces are referred to as ground control stations (GCS) [61]; examples of which include QGroundControl [49], APM Planner2 [51], EOD Workspace by GreenSea IQ [52], NaviSuite Mobula [56], etc. For smartphones and tablet devices, a few applications are available including QGroundControl, DroidPlanner [60], SidePilot [62], and MAVPilot [63]. However, certain interfaces (e.g., SidePilot and MAVPilot) are tailored exclusively for aerial vehicles, relying on GPS signals that are unavailable underwater. Table 1 compares some key aspects and available features of SOTA telerobotic interfaces.
QGroundControl serves as the official interface for submersible (underwater) vehicles within the ArduPilot flight controller developer community [64]. It is built on ArduSub [65], a comprehensive open-source framework specifically designed for underwater vehicles and integrated into the larger ArduPilot ecosystem. Among proprietary interfaces, leading ROV manufacturer VideoRay offers VideoRay Cockpit [53], exclusively designed to operate their ROVs. NaviSuite Mobula is a third-party interface built for VideoRay ROVs and offers some autonomous capabilities including scanning a wall, orbiting a target, etc. Abyssal Offshore [57] integrates real-time XR visualization, $270\degree270$ ° FOV from multiple cameras, and risk analyses for commercial ROV operation. Open-source platforms such as QGroundControl are more suitable for researchers since they allow customization to suit specific hardware configurations and operational needs. Moreover, OpenROV cockpit [66] offers browser-based connectivity without requiring any software installation. Industry-grade interfaces such as FMC Schilling [67] and Teledyne SeaBotix [54] come as rugged hardware packages and do not require external computers or controllers for teleoperation.
4 2.2   Vision-based Consoles
Underwater vision is a well-explored area in telerobotics literature and has long been the key navigational information for human teleoperators [68]. The environmental challenges as well as the inherent limitations of 2D visual systems have been studied for decades and many innovative solutions have been proposed. We summarize the following challenges and corresponding countermeasures employed in vision-based teleoperation technologies. First, traditional consoles with a fixed POV provide limited information to the operator. Monocular FPV cameras lack depth perception and fail to provide peripheral vision, resulting in significantly reduced situational awareness. Second, low-light turbid water conditions reduce visibility and hinder the capture of clear images. Third, artificial lights used on ROVs to enhance visibility have adverse effects too; their bright light gets reflected and back-scattered by suspended particles directly at the front camera, creating glare and large blind spots for the operator [69].
To compensate for poor visibility and to provide the operator a comprehensive understanding of the scene, modern marine telerobotic systems use high definition wide angle cameras [70], structured lighting systems [71], image enhancement [72], and scene parsing technologies [73]. However, interpreting wide-angle fisheye camera images on 2D displays is difficult because of the distortion effects they introduce. Instead, contemporary works present third person view generation process using augmented reality [14], follower ROV [74], external camera [75], etc. Although not popular yet, multispectral and hyperspectral imaging has the potential to provide more robust vision systems with enhanced marine life detection and mapping mechanisms [76]. While vision will remain the fundamental modality for teleoperation, other sensory augmentations are essential for the immersive realization of the remote environment [13].
Besides underwater ROV’s front camera feed, the visual consoles display a wide range of internal and external sensor readings i.e., telemetry data. External sensors capture information such as pressure, depth, temperature, water current, acoustic scans, etc. Researchers emphasize scanning sonar systems (e.g. side scan sonar, synthetic aperture sonar) for safer navigation, as they detect and display obstacles beyond the limited FOV of cameras [77, 78]. Internal data encompasses the status of onboard electrical and mechanical systems including energy consumption, housing temperature, data recording, and streaming status, as well as the ROV’s pose and velocity. Among these, the attitude and heading reference system (AHRS) estimates and displays the robot’s pose on the console which is crucial for maintaining heading and orientation in feature-deprived underwater scenes [79].
5 2.3   Integration of XR
Figure 3: Snapshots of two telerobotics interfaces are shown; (a) An underwater ROV is being maneuvered with human head movements via HMD and haptic feedback controllers; (b) an AR interface with viewpoint augmentation capability [14] is rendering third-person EOB (eye on the back) views [7] for more peripheral information and situational awareness.
Virtually experiencing the depths of the ocean provides humans with an immersive experience, enhancing their perception and enabling more effective engagement with remote underwater settings. In $19931993$, NASA launched the first successful Telepresence Remotely Operated Vehicle (TROV) [46] under the Ross Sea ice near the McMurdo Science Station in Antarctica. This project demonstrated the feasibility of streaming video feeds from ROV and rendering computer-generated graphics on an HMD. Preliminary results also showcased ROV control mechanisms via head movements. However, the progress of XR has been slow in the subsea telerobotics domain, unlike its rapid growth in the gaming industry, surgical robotics, and aerial robotics. According to Zhou et al. [80], an XR model should be able to collect, process, transfer, and reconstruct the immersive scene model of the workspace in real-time and enable intuitive robot controls accordingly. The cost and logistics associated with long-term underwater missions and the unique environmental complexities pose challenges to achieving these criteria [81].
Researchers have come up with innovative solutions to address these challenges and have been successful to some extent in XR interfacing between humans and marine robots. DOLPHYN [82], a mixed reality-based aquatic interface for multi-sensory exploration of scuba diving sites, allows a diver to remotely operate an ROV while being aware of own diving trajectory. Multiple contemporary works explore XR-based view synthesis techniques to provide operators with an enhanced peripheral view during teleoperation. For instance, the augmented visualization of the CoMAS project [83] supports ROV pilots during underwater archaeological site exploration and artifacts collection. A similar approach is presented in VENUS project [84], where the archaeological contents are reconstructed in mixed reality by performing bathymetric and photogrammetric surveys on the real site. Moreover, Ego-to-Exo [14] demonstrate the utility of augmented third-person visuals to increase situational awareness during underwater cave exploration; see Fig. 3 (b). In addition to rendering the visuals, an HMD is capable of actively controlling the ROV [39], freeing the operator’s hands for other navigational or manipulation tasks; see Fig. 3 (a). Windowed or stereoscopic VR [46] further improves perception, when compared against monoscopic VR or desktop-based visuals [85]. Other VR-based frameworks such as TWINBOT [86, 87] offer full teleoperation experiences in simulation, which is particularly useful for pilot training prior to real-world missions. The impact of training and XR visuals is supported by studies showing a reduction in task completion time by over 50%, even when the control method remains unchanged [85]. Two key components of XR systems: simulators and digital twins – are further discussed in Sec. 3.
6 2.4   Multimodalities Interface
Besides vision, human sensorimotor control relies on multimodal sensory feedback such as visual, auditory, and somatosensory (tactile and proprioceptive) cues to interpret the consequences of initiated actions [88, 89, 90]. Haptic feedback is the most recognized one in enhancing XR systems; researchers have explored innovative methods for simulating tactile sensations and haptic devices are employed to deliver haptotactile stimulation (e.g., vibrations and force feedback) synchronized with real-time events [91, 92]. Haptic-integrated interfaces typically feature a visual feed via a desktop or a head mounted display (HMD) and one/multiple haptic feedback devices such as wearable gloves, exoskeletons, body suits, etc. Preferred choices for haptic feedback include vibrotactile, skin indent, pressure, etc. on finger and wrist regions as these options only require cheap hardware compared to exoskeletons or full-body haptic suits. Fig. 3 shows a pair of hand-held controllers for vibrotactile feedback. Other haptic feedback mechanisms and their suitable applications are categorized in Table 2.
The benefits of haptic feedback have been extensively studied for ROV driving and manipulation tasks [93, 94]. Ryden et al. [2] present a framework for assisted manipulation (opening and closing of a valve) using real-time haptic feedback to the operator’s hand. It guides the robotic arm to stay away from forbidden regions using force feedback and pulls the arm toward the desired orientation. Such force feedback is crucial for the operator’s sense of virtual presence and is driven by different sources such as artificial muscle [95], magneto-rheological fluid [96, 97], and dc motors [38]. Researchers also explore how other external sensory information can be incorporated into the feedback system. For instance, Amemiya et al. [98] develop a system that combines water pressure and torsion forces to enhance kinesthetic perception. Moreover, the haptic interface developed by Xia et al. [13, 99] includes the feedback of hydrodynamic state as a vector field in addition to other vibrotactile cues, helping the operator to maintain orientation and heading of the ROV in turbulent water.
Furthermore, whole body motion mapping frameworks [42, 100] take advantage of body dexterity and movements to control multiple motions of an ROV. However, as mentioned in Table 2, haptic feedback on multiple sensory organs may reduce user comfort [101]. To address this, researchers include vestibular feedback using an actuated chair that mitigates HMD-induced motion sickness during long-term operations [42]. Additionally, guiding systems are developed for haptic-supported manipulation that evaluate discomfort in human hand using the Rapid Upper Limb Assessment (RULA) [102], and direct the hand toward a more comfortable posture [103]. Overall, despite the limitations of current haptic devices and the low ergonomics in some applications, the combination of visual and haptic feedback have been demonstrated significant potential to enhance the intuitive sense of telepresence.
Table 2: Haptics and XR capabilities of telerobotic interfaces are categorized and their use cases are highlighted.
  Feedback Feedback Control Ideal Use Cases User Comfort Selected
Location Type Mechanism & Ergonomics References
  Fingertips Vibrotactile Wearable gloves, Grasping, collecting sample, High [37, 2]
stylus controlling valve/knob [104, 105]
Wrist, Force Wearable gloves, Sensing drag and tension, Medium [47, 106, 107]
forearm exoskeleton picking object [108, 109]
Eye Real/virtual HMD Inspecting infrastructure, Medium [46, 85, 86]
video stream long-term exploration [82, 110, 99]
Upper body Tactile, force, Exoskeleton, Sensing and mitigating drift, Low [110, 99, 109]
fluid flow suit sensing velocity and current
Full body Tactile, vestibular, Full body suit, Sensing velocity, pressure, current, Low [42, 111, 100]
force, fluid flow actuated chair using dexterity for manipulation
7 2.5  Integration of Natural Language
NL-based interfaces leverage the natural human way of communicating and reasoning for more intuitive mission planning and control. The process began as experimental research almost two decades ago [112, 113], aiming to simplify human-robot communication by enabling operators to issue high-level commands using everyday speech rather than complex control inputs. Classic visual language systems such as RoboChatTag [114] and RoboChatGest [115] use sequences of symbolic patterns to convey simple instructions from diver to robot/operator, utilizing AR-Tag markers and hand gestures, respectively. Recently, natural language processing (NLP) with deep learning networks has allowed the smooth conversion of spoken/written language into understandable and actionable commands for robotic systems [116]. Early NL interfaces focus on basic navigation commands or status queries such as “take a left", “report battery status" [117], and have since evolved to support more sophisticated mission planning and contextual understanding. Fig. 4 provides a chronological overview of NL integration in telerobotic control and mission management.
Figure 4: A timeline progression of natural language integration in subsea telerobotic interfaces and mission management tasks over the past $1515$ years is shown. The key works referenced are: Hallin et al. [113], PUMMA [118], REGIME [119], Robb et al. [40], RoboChat [116], OceanChat [120], ChatSim [121], OceanGPT [122], OceanPlan [48], and Word2Wave [43].
More recently, the success of large language models (LLMs) has led roboticists to utilize detailed language descriptions to plan robotic missions. LLM-driven mission planning systems reported in contemporary literature are OceanChat [120], Word2Wave [43], and OceanPlan [48]. These interfaces operate as offline planners, processing full mission descriptions and planning the optimum trajectory before deployment [118]. The planner first converts spoken or written instructions into structured commands using speech recognition and syntactic parsing. This is followed by semantic analysis, where the system interprets the meaning and maps it to specific robotic actions or a sequence of tasks. A particular example for Word2Wave interface [43] is shown in Fig. 5. When an operator specifies: “spiral for $55$ turns to a radius of $5050$ m at a speed of $11$ m/s”; the system breaks down the request into individual sub-tasks and plans accordingly such as setting velocity, setting altitude, and finally making spiral move to reach the specified radius. Such interfaces enable a robotic system to plan and execute complex missions from high-level operator instructions.
In contrast to offline planning interfaces, online interfaces allow real-time dialogue between the operator and
Figure 5: An example of a language-guided mission planning interface [43]. The robot parses verbal instructions into sub-tasks such as S (start), Mv (move), Sp (spiral), Az (azimuth), E (end); and then generates mission waypoints.
the robot. Hastie et al. [119] demonstrate an online dialogue module named REGIME (Report Generation from Metadata), where sensor data and event logs are summarized into mission updates upon operator demand. Robb et al. [40] introduce a chat-based interface for controlling a group of robots, allowing the operator to inquire about the status of individual robots via the leader and adjust their behavior as needed. This reduces the cognitive load on operators by shifting focus from managing multiple robots to a single control point through the leader. In ChatSim [121], dialogue capabilities are utilized to customize a simulated environment. The operator first asks for specific structures (e.g. canyon, shipwreck, coral reef, oyster) to be added into the virtual world, followed by issuing commands to the robot such as “go through the canyon and find the shipwreck”, “move in a circle of radius $33$ m over the oysters”, etc.
The current challenges with integrating language models into telerobotic systems are multifaceted. First, the computational demands of LLMs may exceed the capabilities of resource-constrained underwater systems. To address this, researchers are developing smaller language models (SLMs) while technologies such as LoRA [123] are helping to adapt LLMs for real-time applications in edge devices. Second, ensuring that the robot accurately interprets commands and understands spatial references requires sophisticated models, extensive training data, and robust error-handling mechanisms [116]. The consequences of misinterpretation can be severe in high-risk subsea operations, hence current research focuses on making these systems more resilient, context-aware, and capable of handling complex, multi-step instructions [124]. Moreover, online commands can be error-prone due to variations in dialect, tone, or pronunciation [125]. Instead, an offline planning interface allows the operator to review and correct for potential errors before launch. Further discussions on this are presented in Sec. 6.
Figure 6: A digital twin of an underwater ROV is shown in action; (a) a human is teleoperating it to pick up an object using head movements and finger grips; (b) the twin model rendered in HoloOcean is mimicking the action in virtual environment.
8 3 Simulators and Digital Twins
Simulation environments offer a valuable alternative for prototyping, error debugging, and pilot training for marine robotic systems in a controlled, risk-free setting, without requiring expensive logistics and access to real waterbodies. These simulators not only replicate the dynamics of underwater conditions but also support the integration of digital replicas of physical systems, facilitating the testing and development of robust control algorithms and mission planning strategies for real robots. Fig. 6 shows an example of underwater grasping with a digital twin rendered in the HoloOcean simulator.
9 3.1  Digital Twins (DTs)
A digital twin refers to a simulated replica of a real robot that possesses similar attributes and can mimic the motion of its real-world pair within a virtual environment [126, 127]. The concept of DT for field robots was first presented by NASA, where the digital entity was defined as an integrated multi-physics, multi-scale, probabilistic simulation of a vehicle or system that mirrors the life of its real-world peer [128]. This project envisioned four major applications of the DT technology.
  • Pre-mission simulation: The twin will simulate the vehicle’s future missions before launch, enabling studies of mission parameters, anomalies, and mitigation strategies to enhance mission success.
  • Real-time flight mirroring: The twin will continuously update and mirror the actual flight with real-time sensor data, predicting flight status for ongoing analysis and adjustment.
  • In-situ forensics: The twin will allow diagnosis and analysis of potential catastrophic faults in real-time and address anomalies during the flight.
  • Mission modification analysis: The twin will assess the impact of unforeseen mission parameter changes (e.g., failed actuators), and analyze new load distributions to inform mission decisions.
DTs exhibit varying levels of complexity, ranging from basic motion replication to comprehensive emulation of physical and operational attributes. The simplified setup shown in Fig. 6 does not replicate the real testbed in the simulator; however, the digital robot mimics the motion of the real ROV, thereby qualifying as its twin. For more sophisticated DTs, the digital replica must mirror key physical properties (e.g. dimensions, weight, buoyancy), energy consumption profile, sensor configurations, and interaction with the external environment. Examples of some SOTA simulators and their DT collections include BlueROV2 in UNav-Sim [129], Triton XLX in PaleBlue [130] and VMAX simulator [131], VideoRay Pro5 and Defender in Era’s VideoRay simulator [132], and Girona500 in UWSim [133]; see Fig. 7.
Figure 7: Snapshots of a few marine robotic simulators are shown; the top row shows proprietary simulators for professional use [130, 134, 131]; the bottom row shows open source simulators mainly used for academic research [26, 27, 133].
The use of DTs has been reported in different branches of robotics including aerial, manufacturing, surgical, space, and marine robots. All DT systems are built upon some common backends with domain-specific techniques for data acquisition and rendering [127]. Generally, the raw sensor data acquired from the robot is processed with physics engines and reconstructed as a high-fidelity DT model [127]. Considering the complex nature of subsea environments, Van et al. [32] identify the communication between the two entities as a fundamental component of a successful subsea DT system. The authors highlight three use cases for DT: (i) ROV launch and recovery training, (ii) failure prediction and remaining useful life estimation, and (iii) trust-aware DT development to reduce operator’s decision-making time in critical situations. However, other literature [30, 135] argue that DT systems do not require a fully reliable or accelerated communication mode. Instead, during a communication disruption, the predictive nature of a DT system can be utilized to create a realistic representation as if communication had remained uninterrupted. DT systems have also been deployed in underwater sensor networks to test communication reliability and synchronization accuracy between multiple nodes [136]. More recently, Adetunji et al. [29] integrate SLAM and path planning capabilities in a versatile DT system. Other applications of DT in marine robotics include pipeline inspection [137], assistive control [13], trajectory planning [30], etc.
10 3.2  Underwater Simulators
A simulator is essential to work with DTs as well as to mimic the external environment parameters. However, given the complex fluid dynamics and other physicochemical variations, it is extremely difficult to design an accurate physics-based model of an underwater environment. Engineers have created aquatic environments with physics and game engines (e.g. Unity, Blender, WebGL3D, ODE) as backends and have integrated basic properties of the medium such as buoyancy, waves, current profile, etc.; examples include UUV Simulator [26] and UWSim [133] that support Gazebo plugin [138] as well. Independently designed virtual ROVs or DTs of real ROVs are rendered in these platforms to test their kinematic properties and interactions with the surroundings. Researchers have studied different underwater vehicles such as gliders [30, 139, 140], ROVs [135, 26], AUVs [137], underwater legged robots [141], and humanoid robots [111] in these simulators. Yang et al. [30] investigate a glider trajectory in a deep virtual ocean. Their simulations show the effects of the current gradient across the water column on the glider’s trajectory, especially during the vertical dives. Hu et al. [139] model a blended-wing-body underwater glider (BWBUG) and investigate the accuracy of a pitch controller under the influence of ocean currents.
Table 3: Comparison of underwater simulators are shown. Basic state estimation sensor plugins (e.g., IMU, GPS, depth) are common for all simulators, therefore omitted here; indicates digital twin of a real ROV.
(a) Comparison based on sensor integration capability, scenario libraries, and available robotic agents.
  Sensor Plugin Built-in Scenarios Robot Models
  PaleBlue [130] RGB-D camera, laser Reefs, oil and gas rigs Triton XLX
Abyssal [134] Camera Rigs, sand cloud, pipelines Abyssal ROV
M8-R Sim [142] Camera 140+ coral reefs with starfishes M8-R ROV
Era VideoRay [132] Camera, sonar, lights 6 seabed scenarios with shipwrecks VideoRay Pro5 , Defender
GRi VROV [67] Camera and sonar Pipelines and rigs LittleGeek, BigGeek
UUV Simulator [26] Camera, sonar, DVL, CPC Lakes, shipwrecks, coast, BOP panels A9 , LAUV , Saga , RexROV
UUVSim [143] Camera, sonar Open water, current, seabed 6T UUV, RexROV
HoloOcean [27] Camera, sonar, beacon, DVL Dams, pier harbors, pipelines, open water 6 ROV agents
UWSim [133] Camera, DVL, multi-beam sonar Dredging, pipelines, seabed with amphora Girona500
UNav-Sim [129] Camera, LiDAR Seabed with algae and sand-covered pipelines BlueROV2 heavy
MARUS [144] Camera, LiDAR, sonar, DVL, AIS Ships, human divers, seabed One AUV
StoneFish [145] RGB-D camera, sonar, DVL Sea surface, seabed, wind, sun Girona500
NaviSuite [56] Camera, sonar, DVL, USBL Cables, pipelines, and ship anchors VideoRay Defender
VMAX Simulator [131] Camera, sonar Dynamic seabed, cables, ship with crane Triton , Comanche , Mojave
ACSM [146] Camera Rigs and pipelines Triton XLX
Unicom [147] FPV and TPV cameras Pipelines and rigs with valves, knobs, doors 2 work class ROVs
(b) Comparison based on the availability of various features and supported platforms.
  Pale Abyssal M8-R Era GRi UUV Holo UW Unav MARUS Stone Navi VMAX ACSM Unicom
Blue Sim Sim VideoRay VROV Simulator Ocean Sim Sim Fish Suite Simulator
  Open Source
Haptic Support
ROS Support
Training Modes
Dynamic Obstacles
OS (W,M,L) ✓✕✕ ✓✕✕ ✓✕✓ ✓✓✓ ✓✕✕ ✓✓✓ ✓✕✓ ✓✓✓ ✓✕✓ ✓✕✓ ✓✓✓ ✓✕✕ ✓✕✕ ✓✕✕ ✓✓✓
Fig. 7 illustrates some SOTA underwater simulators developed across academia and industry. Among proprietary simulators, Abyssal [134] and ACSM [146] offer a rich collection of underwater infrastructures and work-class ROV models, designed to train operators in tasks such as underwater mining, pipeline repairing, and maintenance of offshore oil and gas rigs. PaleBlue [130] and VMAX Simulator [131] allow users to import custom-designed ROVs from CAD models and simulate ROV behavior, forces, frictions, etc. in the virtual world. Era Marine and EIVA have partnered with VideoRay to create advanced simulators for the VideoRay Pro5 and Defender ROVs. These simulators, Era’s VideoRay [132] and EIVA NaviSuite [56], provide advanced tools for planning, surveying, and underwater inspections. M8-R Sim [142], originally developed for the MATE ROV Competition, has over $140140$ simulation scenarios, including those for monitoring and estimating the growth of Crown of Thorns Starfish (COTS). Unlike higher-end professional simulators, open-source alternatives offer fewer built-in scenarios and ROV models but allow for better customization and further development according to mission-specific needs. For instance, UWSim [133], StoneFish [145], and HoloOcean [27] support numerous sensors e.g., side scan sonar, single beam and multi-beam imaging sonar, acoustic beacon, DVL, magnetometer, force sensor, etc. UUV Simulator [26] integrates chemical particles and a simulation scenario to measure their concentration using a CPC (condensation particle counter) sensor. They also provide closed water (lake) and open water (coast) scenarios, each with different current profiles. As shown in Table 3, the open-source solutions are better suited for research and development with ROS support, whereas licensed products offer superior training opportunities for commercial ROV operation.
11 4 Human-Machine Shared Autonomy
Telerobotic systems are categorized into three types according to their intelligent decision-making capabilities: (i) direct teleoperator control with no intelligence, (ii) shared/collaborative reasoning, and (iii) autonomous/self-sufficient systems. In direct teleoperation, the operator has total control over the robot’s motion and actions, generally through a live feed and a motion controller. However, these methods are constrained by communication latency and operator workload, making them less efficient for long-term complex tasks. Shared autonomy features bridge the gap between the human and the machine by allowing robots to perform certain tasks autonomously while the operator manages high-level decisions [35, 22]. This reduces the operator’s cognitive load, compensates for communication delays, and enhances operation safety. In contrast, AUVs function independently with minimal human intervention [148]. These robots rely on advanced algorithms for perception, decision-making, and execution, making them ideal for repetitive tasks in familiar environments. However, their utility is constrained by environmental uncertainties, the cost of logistics, and the limitations of current AI-based technologies [149, 150].
12 4.1  Direct Teleoperation Mode
Direct or hands-on teleoperation involves controlling the movements of an underwater robot using low-level commands without intelligent decision-making by the robot. The advantage of such systems is that the operator can execute instant decisions during the operation. It allows for responsive control, enabling precise adjustments that automated systems may struggle with. Despite high cognitive demand, direct teleoperation remains the preferred approach for manipulation tasks involving high risk and fine precision, where the operator controls one joint angle at a time in a joint-by-joint fashion using a traditional console [151]; examples include wreck inspection [152], valve control [153], benthic sampling [151], etc.
These systems are challenged by the inherent communication delays in underwater environments, which can reduce control precision and responsiveness. The delays further complicate tasks that depend on real-time feedback [154, 153]. To address this, researchers have developed adaptive control strategies to compensate for both model uncertainties and external disturbances during teleoperation [155, 153]. Furthermore, electromagnetic and optical communication modes offer promising solutions for establishing high-speed reliable links [156, 157], improving the real-time commands and feedback flow in direct teleoperation systems.
Table 4: The table summarizes various types of intelligent machines and their autonomous task-sharing capabilities reported across the subsea telerobotics literature.
  Type of Human (Teleoperator) Side Machine (Underwater ROV) Side Selected
Intelligence Responsibility Cognitive Load Intelligent Capabilities Major Applications References
  Command Issue high-level Medium Translate to low-level action Surveying, grasping, [158, 34]
interpretation commands and check feasibility inspections, interventions
Online course Set new goal point Low Update local planner and Dynamic path planning [20, 48]
adjustment or trajectory mission parameters in uncertain environment
Error Supervise and act Medium Diagnose issues, report, Fault tolerant system for [159, 160]
handling upon anomaly attempt resolution long-term exploration
Multi-robot Operate the Medium Coordinate and control Coverage, inspection, [161, 9]
collaboration leader ROV follower ROVs search and rescue [6, 162]
Mission Set mission Low Generate waypoints, Exploration, mapping, [120, 48]
planning objectives execute motion area coverage [30]
Learning from Demonstrate Initially high Learn knowledge base, Routine maintenance, [163, 164]
demonstration relevant tasks adapt to novel tasks sample collection [104, 165]
Context-aware Provide contextual Medium Search for relevant features, Inspection, surveillance, [5, 166]
exploration cues or guidance report key findings environmental monitoring [167]
Docking Guide the robot Low Identify docking point, Charging, data transfer, [168, 169]
assistance close to the dock self-correct fine motion repair, recovery [170, 171]
13 4.2   Human-Robot Collaborative Reasoning
The goal of collaborative decision-making systems is to offer a blend of human oversight along with automated functions. These systems allow operators to set high-level tasks or objectives while the robot handles low-level operations, thus reducing cognitive load on the operator and enabling more efficient operation [22]. For instance, an operator can command the simple autonomous task of hovering; the robot determines its state using an IMU, DVL, and depth sensors to correct for any drift motion to remain stationary. More advanced reasonings are achieved in a supervisory manner (for a single robot) or with a leader-follower strategy (for a group of robots). Table 4 provides an overview of the SOTA intelligent capabilities of underwater robots and their task-sharing mechanisms with humans.
Mission planning, online mission adjustment, and fault diagnosis systems benefit from shared autonomy while keeping the operator in the loop [172], but without demanding continuous oversight. Research in supervised planning and navigation demonstrates that a new mission path suggested by a circumstance-aware robot is often safer and reduces task completion time [34, 20, 48, 173]. The GALILEO planning interface adjusts the course of actions and resource allocations based on high-level goals set by the operator [174]. Additionally, intelligent robots are capable of detecting and reporting internal faults [175] and data anomalies [159], reducing the need for constant operator oversight. Such distributed control also allows the operator to focus on mission objectives while offloading low-level tasks to the robot. Sharing scientific objectives with a context-aware robot is beneficial too, as the robot can search for specific cues or features under operator guidance. The CView project [5] provides an example; the operator defines the object of interest (e.g., wall, vessel, sluice) and the features to detect (e.g., cracks), enabling the robot to survey the objects and report key findings.
SUPERMAN, the first supervisory subsea teleoperation interface, was proposed by NASA in $19781978$ [176]; however, the development has been slow due to the difficulty of designing high-level autonomy considering the environmental challenges. SAUVIM (1998199819981998[177] is one of the earliest semi-autonomous vehicles with a 6-DOF manipulator arm developed for deep water intervention tasks. More recently, supervisory control frameworks such as DexROV [109] and SHARC [21] have been developed where the operator performs remote manipulation tasks using language commands and hand gestures while experiencing the task in immersive virtual reality [34]. The GRASPER project [178] utilizes a pre-scanned scene point cloud to inform the robot about obstacle locations. The autonomous action is rendered in the UWSim simulator [133] in real-time, allowing the user to supervise grasping tasks in a virtual world.
Moreover, contemporary researchers have developed adaptive supervision systems where the robot understands the operator’s overarching intention and adjusts the low-level reactions accordingly. Lee et al. [20] propose a shared teleoperation method where humans can provide real-time input by sketching a desired path on the visual console, guiding the intelligent system, which plans an optimum path along that drawn trajectory and executes the motion. The adaptive supervision system built within the GreenSea interface provides operators customized assistance based on their skill levels [35]. This helps novice pilots achieve performance comparable to that of experts.
For more complex scenario, learning from/by demonstration (LfD/LbD) represents another collaborative paradigm, where the robot learns a knowledge base from operator-provided demonstrations and uses it to autonomously perform analogous tasks in the future. Unlike reward-driven reinforcement learning (RL), LfD relies on expert demonstration, resulting in behavior that closely mirrors the operator’s approach [179]. LfD has been extensively applied in subsea manipulation tasks, particularly for drilling and cutting [180], valve control [181], hot-stabbing [182], etc. A major limitation of many LfD approaches is they require optimal demonstration to learn, which is not always feasible, particularly for path planning. This is addressed by accounting for the variability in demonstration quality and combining Bayesian learning with probabilistic path planning [165]. Another challenge of LfD is its struggle to resolve discrepancies between the demonstration space and the robot’s operational space. To address this, Havoutis et al. [163] propose a probabilistic framework utilizing a hidden semi-Markov model to construct robust task representations. Their approach achieves a ten-fold reduction in manipulation task completion time using a simulated DexROV; however, they did not consider vehicle dynamics or external disturbances encountered in real-world environments. Among real-world projects, PANDORA [164] demonstrates the role of LfD in automating inspection and intervention tasks at subsea oil and gas facilities. The authors emphasize the necessity of a cognitive knowledge base enriched with semantic information about the workspace, as opposed to relying solely on pre-programmed trajectories or task-specific descriptions.
In multi-robot systems, a directly teleoperated robot, known as the leader, works alongside one or more follower robots [173]. These followers possess some local autonomy and operate as a coordinated group. The human operator, controlling the leader, oversees the collective movement of the fleet [183]. Two fundamental challenges in such fleets are formation control and trajectory control [184]. The operator plays a crucial role in modulating the fleet’s velocity through the leader to ensure the followers maintain formation, avoid collisions, and stay synchronized with the leader. Such robotic fleets, also known as swarms, are useful for tasks requiring extended area coverage, dense scans, and rapid completion. Notable applications include search and rescue [161], structure inspection [6, 185], ship hull inspection [186, 187], large area mapping [9, 188, 189], cooperative localization [190], and continuous surveillance of underwater facilities [162]. The complexity of these tasks exceeds the capabilities of a single vehicle or imposes an excessive mental load on the operator. In contrast, a leader-follower multi-vehicle configuration can accelerate task completion, enhance vehicle cognition, and reduce the operator’s load.
A recent survey by Zhou et al. [191] characterizes operators as independent from the swarm, limiting their role to mission initialization and maintenance; however, this perspective does not cover the full spectrum of multi-robot literature. For instance, several studies [192, 193] explore scenarios where follower robots autonomously estimate their relative pose while the operator directly controls the leader. Moreover, contemporary research emphasizes heterogeneous robot groups, unlike traditional homogeneous swarms, where all robots had the same characteristics and role [194]. In heterogeneous settings, the operator allocates tasks based on each robot’s specific capabilities and real-time operational status [195]. For marine applications, heterogeneous groups typically include an unmanned surface vehicle (USV) that acts as a communication relay between ROVs and the operator; an example is demonstrated in [6], consisting of one ASV and two SWIM-R ROVs for shallow water pipeline inspection. These configurations enhance mission reliability by leveraging the USV as an additional wireless communication channel.
14 4.3   Autonomous Systems
As opposed to direct teleoperation and shared autonomy schemes discussed above, autonomous systems or AUVs operate independently, requiring no control commands from the operator. Subsequently, HMI for an AUV is primarily designed for status updates and mission management rather than direct control. However, even fully autonomous systems rely on human operators for mission planning, progress monitoring, and emergency intervention. This highlights the critical role of AUV interface design in ensuring the safety and effectiveness of autonomous missions.
Generally, the interface includes a software console, a surface control module (SCM), and a transducer to maintain real-time communication with the AUV. Available real-time commands at the operator end are typically limited to starting or aborting the mission, requesting status updates, or intervening during critical moments. Contemporary works have integrated more advanced automated features into these interfaces. For instance, the MARES AUV interface automatically performs pre-mission procedures such as estimating mission duration and path length, testing communications, etc. [196]. Kaeli et al. [159] develop an anomaly detection framework that enables AUVs to focus on high-resolution scanning and transmit only the anomalous data to the operator for inspection, reducing data congestion and improving operational efficiency. A similar approach is reported for crack detection in subsea pipelines using a camera, laser, and multi-beam echosounder (MBES); the AUV detects potential cracks and sends the scanned images to an operator for further review [5]. The operator then requests close-up detailed examination if necessary, or continues the mission.
Commercial AUV/UUV Platforms. Currently, the top application segments for AUV/UUVs are military and defense, offshore oil and gas drilling, oceanography, archaeology, and alternative energy harvesting [150]. Their operational costs are high due to factors including maintenance, repair, data acquisition, personnel training, and mission-specific logistics. Consequently, access to AUVs has been restricted to governments and their agencies, industry leaders, and a handful of academic and research labs. Examples of SOTA AUVs are L3Harris IVER (US), A9-E ECA (France), Teledyne Gavia (US), Kongsberg Hugin (Germany), SAAB Seaeye Sabertooth (Sweden), Exail A18 (France), and Atlas Elektronik SeaCat (Germany). Another lightweight low-cost category known as micro-AUV is gaining popularity for academic and research purposes; examples include NemoSens (France), Seaber Yuco series (France), and EcoSUB (UK). A detailed review of current AUV technologies, their challenges, and applications is outlined in [197].
15 5  Engineering Integration of Interactive Control
Figure 8: Examples of bidirectional human-robot interaction in telerobotic tasks are shown, highlighting modern control and feedback mechanisms. To this end, a few prominent research topics based on task domain interaction are: manipulation: [2, 104, 93, 198, 163, 83], mission planning: [118, 43, 30, 120, 122], navigation: [79, 110, 199]; and application domain interaction are: inspection: [200, 164, 5, 186, 6], search and rescue: [201, 161], and mapping: [28, 9].
An end-to-end telerobotic system is realized as a loop of perception, sensory feedback, and control commands. The robot first perceives the environment using multiple sensors, determines its state, and transmits the information to a surface station. The operator interprets the remote information through various feedback mechanisms such as visual displays, auditory cues, and haptic sensations. Based on the feedback, the operator uses his cognitive intelligence to plan the motion and send control commands to the robot via switches, GUI buttons, gestures, verbal language, etc. Maintaining this loop of action is crucial for ensuring safe navigation and task completion. Fig. 8 demonstrates some examples of human-robot interaction through command and feedback exchanges during teleoperated missions.
16 5.1  Visual Feedback Technique
Visual feedback is typically sent from cameras mounted inside a pressure-sealed housing, providing a front-facing first-person view of the robot. The camera sensor setup and lens type can vary widely; monocular and stereo cameras with optional IR sensors are common choices since they can provide low-light vision and depth perception. The lens can provide a wider FOV, as high as 220220220220°, representing a fish eye lens [75] that offers panoramic view of the surroundings. Traditionally, the camera feed is viewed on the surface computer’s screen from a fixed viewpoint; however, recent technologies such as camera arrays, HMD, and view synthesis algorithms offer augmented peripheral view to the operator.
HMD. In addition to displaying real-time video feed, HMDs also facilitate multi-sensor data integration, allowing operators to simultaneously monitor environmental parameters, sonar scans, and other telemetry data within the display. HMDs become even more effective when paired with a movable camera setup on the ROV [202]. As the operator moves their head, the ROV’s camera mimics these movements, providing a natural and intuitive way to observe the surroundings.
Augmented Viewpoints. Unlike the limited first-person view from the ROV’s front camera, augmented third-person perspectives offer a chase view [203] or an overhead view, enhancing spatial awareness. By toggling between first- and third-person perspectives, operators can transition from detailed, close-up control to a broader peripheral view, reducing the risk of collisions. Researchers have explored both fixed [204, 199, 7] and dynamic [205, 206, 14] viewpoint augmentation techniques, the latter being computationally expensive but providing a continuous mosaic-like view of the periphery. Augmented third-person views, when combined with real-time data overlays such as sonar or lidar maps [199, 207], further render a rich virtual environment for interaction.
17 5.2   Haptic Feedback Technique
Haptic feedback involves sensory information derived from mechanoreceptors embedded in the skin (tactile input) and muscles, tendons, and joints (kinesthetic inputs), provided to the human operator [208]. Haptic feedback in ROV teleoperation is crucial for enhancing the operator’s sense of control and situational awareness, improving task performance, and reducing cognitive load during complex missions. The existing literature acknowledges the extremity of underwater environments and the challenges of maneuvering an ROV using a video stream alone. As such, haptic feedback is augmented into HMIs for improved control [80]. They are primarily used to warn of potential collisions and sense the ROV’s interaction with the environment [2]. Haptic interactions are broadly categorized into two types: kinesthetic (which involves force) and tactile (which involves touch and vibration).
Tactile Feedback. Tactile feedback enables humans to recognize local properties of objects being manipulated such as shapes, edges, and textures, based on measures of contact forces on the operator’s skin [209]. Ideal areas for delivering this feedback include the fingertips, palm, and wrist, as they are rich in somatosensory nerves.
Figure 9: A haptic teleop interface with gloves and bodysuit [13] is shown.
Xia et al. [13] use HaptX gloves [210] for a multi-level pneumatic haptic interface design; see Fig. 9. The $130130$ pressure points within the user’s palm and the airflow channels along the fingers make it a suitable glove to sense micro-scale turbulence as well as palm-level vibrations. Their multi-level design allows the operator to choose the appropriate level of feedback based on the task’s complexity. For instance, high-resolution turbulence data is vital for accurate docking and inspection tasks, but excessive feedback during routine maintenance or exploration may overwhelm operators, leading to cognitive overload without improving efficiency [211]. While the feedback is engineered on land, sensing the raw information in deep water adds to the challenge since the sensors become exposed to the high pressure of the water column. Kampmann et al. [93] address this by developing SeeGrip, a rugged multi-limb manipulator capable of providing tactile sensations from deep water. To achieve a higher depth rating, a fiber-optic force sensing mechanism is proposed, unlike capacitive sensors [212] or conductive rubbers [213] commonly used for terrestrial robotic arms.
Kinesthetic Feedback. Kinesthetic feedback provides information on the position, velocity, force, and torque of objects through receptors in muscles and joints [214, 215], allowing operators to perceive the weight of objects or the force required to interact with them. This type of feedback is essential for precision tasks such as grasping delicate objects. Khedr et al. [216] design a gripper jaw that delivers force feedback to the operator’s fingers, mimicking the stiffness of the grasped object. The utility of force feedback has also been explored for sensing orientation and detecting collisions. For instance, an additional hand-held steering mechanism replicating the ROV’s rotation is useful, as demonstrated by Shazali [107] and Abdulov et al. [217]. The manipulator developed by Ryden et al. [2] utilizes non-contact (camera) sensors to identify nearby obstacles and provide force feedback to the stylus controller, helping to prevent unwanted contact or collisions during manipulation. However, waterproofing bulky force sensors in high-pressure environments while maintaining sensitivity remains a significant challenge [218]. Furthermore, the force-generating actuators tend to be large and power-intensive, leading to tactile feedback being the preferred choice for most applications.
18 5.3   Pseudo-Haptic Feedback
Given the high costs and logistical challenges of open water ROV deployment, virtual environments are beneficial for prototyping and user training. In the virtual environment, pseudo-haptic feedback creates a haptic illusion to operator, allowing them to perceive sensations such as friction and drag forces while observing the task from visual feedback [219]. Lecuyer [220] utilizes pseudo-haptic feedback as a means for simulating the sensations of stiffness, texture, and mass of a virtual object through visuo-haptic perception. The virtual platforms can provide operators with realistic subsea sensations [100], which is particularly helpful for novice operators to prepare for real-world missions [99].
Figure 10: A chronological progression of human-robot interaction technologies is shown, focusing on gesture-based and language-based methods. CADDY [221] and ScubaNet [222] are two comprehensive dataset for diver gesture recognition. Other key works referenced are: RoboChatTag [114], RoboChatGest [115], DexROV [109], Pandora [164], Mi et al. [41], Robb et al. [40], Chiarella et al. [223], Fulton et al. [224], Kapicioglu et al. [225], Xia et al. [13], and SHARC-VR [21].
19 5.4   Other Feedback: Logs, Maps, Auditory Cues
In addition to visual and haptic feedback, teleoperators receive information about the robot’s surroundings through other sensors. For instance, rotating sonars and lasers detect obstacles within a certain range and present the data as occupancy maps. Environmental parameters, including depth, temperature, and physicochemical properties (e.g., pH, salinity, oxygen levels), are recorded by specialized sensors and stored as mission logs or time-series data. All this information is typically displayed on a screen, overloading the pilot’s visual channel. Therefore, researchers suggest auditory cues as a means to reduce visual workload during teleoperation. The integration of auditory cues was initially explored through Audio Augmented Reality (ADAR), which combines augmented reality with a virtual audio-visual interface to facilitate improved ROV navigation  [226]. However, the auditory cues used in this work are simply verbal commands (e.g., left, right), making it similar to NL-based interfaces. Subsequently, this approach is expanded with the development of an auditory guidance system [227] that utilizes non-verbal audio (i.e., auditory icons) via sonification technique [228]. These works demonstrate that the concept guidance-by-sound is feasible and more useful in noisy low-light oceanic environments [229]. Nevertheless, the authors point out that auditory feedback has much lower spatial resolution compared to vision, as humans are far more adept at visual navigation [230].
20 5.5   Interactive Control
The existing works explore various mechanisms to issue control commands to ROVs, each with unique strengths depending on the task and the operational environment. Traditional joysticks and graphical user interfaces (GUI) offer a familiar and reliable method, where operators use physical controllers or software dashboards to execute specific actions. Gesture-based control systems, utilizing motion sensors and IMUs, provide an intuitive way to interact with the robot, translating the operator’s hand/body movements into robotic actions [99, 164]. Lastly, natural language (NL)-based controls introduce hands-free natural communication, allowing operators to streamline complex sequential tasks with spoken instructions [43, 21]. Fig. 10 summarizes some prominent research works on enhancing human-robot interaction via gestures and NL-based control modes.
Handheld Controller and GUI. The traditional control method, consisting of a software console and a hardware device (e.g., joystick), remains the primary approach for teleoperating subsea robots to this day. In $19781978$, NASA introduced a comprehensive command and control (C2) interface that combines a joystick, keyboard, specialized button box, and a GUI for undersea manipulation [176]. While these devices offer more direct, low-level control with minimal risk of interpretation errors by the robot, they demand constant attention and keep the operator’s hands engaged. As a result, such systems are only suited for hands-on teleoperation with limited robotic autonomy.
[…]
Figure 11: Snapshots of natural language-based interactions are shown; (a) waypoint mission planning from verbal command [43], status query and real-time written response generation [40], and (c) command to execute a specific task [48].
Language Commands. Verbal and written NL commands represent the latest and most intuitive mode of controlling underwater robots. SOTA research showcases the potential for real-time dialogue between humans and robots, moving beyond pre-programmed commands to more flexible mission management and conversational interfaces; see Fig. 11. Earlier interfaces such as REGIME generate post-mission summary/report from mission logs and sensory data [119]. MIRIAM [235], the successor of REGIME, integrates real-time mission status query options and provides important notifications such as objective completion and fault occurrences. In the Neptune autonomous command and control interface, Robb et al. [40] further demonstrate MIRIAM’s value in multi-robot coordination, where the operator communicates directly with the leader robot that relays information to/from the followers. For instance, when the operator asks “how many survey legs vehicle-2 has completed"; the leader retrieves and relays that information from vehicle-2 to the operator. These interfaces operate on written instructions, hence the voice-to/from-text conversion is not thoroughly tested. More recent works utilize LLMs to address this, as demonstrated in RoboChat [116], OceanChat [120], and OceanPlan [48]. These works highlight the ease of using verbal commands for both pre-mission planning and real-time queries or adjustments.
[…]
21 6 Challenges and Open Problems
The subsea telerobotics domain presents unique challenges due to the hostile and dynamic nature of the underwater environment. Key technical hurdles include limited FOV, latency in long-distance operations, hydrodynamic variations, uncertainty in state estimation, and the presence of dynamic obstacles within confined spaces. Open problems in this sector span the fields of robotics, low-light perception, AI, computer vision, human cognition, and ocean engineering. Fig. 12 highlights the promising directions for future research. Few prominent research areas and their open problems are discussed below.
Figure 12: Key open research problems in the subsea telerobotics are summarized. Operator-centric research emphasizes enhancing telepresence through immersive feedback and effortless control strategies, while ROV-focused research prioritizes advanced sensing capabilities and intelligent decision-making.
22 6.1   Human-Machine Dialogue in Natural Language
As natural language has been proven to be one of the most effortless control methods for teleoperation, the future goal is to enable real-time natural conversations, where the robot would possess a cognitive understanding of the goal and actively engage in decision-making. LLMs, particularly GPT models [237, 238], have successfully conducted real-time human-machine dialogues for various HRI settings [239, 240]. However, marine robots face the additional challenges of integrating sensory data and mission objectives into their decision-making process while executing a teleoperator’s command. ChatSim [121] and OceanGPT [122] have shown some progress in this regard, demonstrating limited dialogue capabilities and gathering vast knowledge of the ocean environment into the model, respectively. Further research can enable the robot to understand various instructions, reason logically, and propose alternative strategies using a cognitive understanding of a mission.
While language models can parse instructions and reason from a long conversation, they struggle with linguistic biases [125] such as variations in dialect/pronunciation, indirect commands, as well as mission-specific terminologies. Therefore, another research direction would be a risk assessment for voice-operated subsea missions and developing fail-safe mechanisms for the robot to identify, query, and correct ambiguous or potentially dangerous commands.
23 6.2   Optimizing Learning from Demonstration (LfD)
Robot’s ability to learn from repetitive tasks and apply that knowledge to unknown scenarios remains under-explored in marine robotics, unlike terrestrial and aerial domains [241]. Few recent works have tested LfD for underwater manipulation [163, 164], and sample collection [165]. However, extensive real-world evaluation is necessary to adopt these developments into industrial applications.
In general, any reinforcement learning (RL) pipelines struggle with dynamic changes such as robot’s component failure or unseen terrain [242]. Extensive training (i.e., demonstrations) in diverse scenarios is a straightforward solution. However, crowdsourcing demonstrations are extremely challenging in harsh subsea environments, unlike ground/aerial scenarios [241]. This motivates multiple research directions. First, the robot should learn how to adapt to an unseen task using prior experience. Meta RL [243] and online RL [242] have shown progress in such model adaptation through reward-driven optimization. Researchers have utilized these approaches for improving subsea locomotion [244], path planning [245], and target tracking [246]. Second, realistic simulators featuring various underwater scenes can be a potential alternative for recording numerous task demonstrations. Third, algorithmic studies are needed to determine the minimum number of demonstrations required to achieve a desired performance threshold. Additionally, hybrid approaches that combine LfD and RL should be explored, where operator interventions during autonomous missions are treated as demonstrations or corrective feedback. Such interventions could potentially teach the robot to generalize knowledge, refine policies, and avoid similar errors in the future.
24 6.3   Hydrodynamic Sensing & Haptic Feedback Modeling
Modeling haptic sensations from remote underwater settings presents multi-faceted challenges. First, the extreme pressure and complex hydrostatic/hydrodynamic forces limit the sensing accuracy [13]. Second, existing fluid dynamic models cannot accurately represent detailed information for different spatiotemporal scales of turbulent water flows. Third, translating high-fidelity hydrodynamic data at both the micro and macro levels in intuitive haptic feedback that effectively engages teleoperators’ sensory perception while maintaining ergonomics – remains as open problem.
In recent years, researchers have explored materials such as Polycrystalline Nickel Titanium [247], Carbon Nanotubes [248], and Tungsten Disulfide Nanosheets [249], which exhibit flexible, hydrophobic, and corrosion-resistant properties. Doppler current profiler and artificial lateral line sensors have shown promising results for far-field and near-field flow measurements, respectively [99]. Other techniques such as Smoothed Particle Hydrodynamics (SPH) [250] and Position Based Dynamics (PBD) [251] are being explored for modeling free surface flows. Potential research directions in this regard are physics-driven learning [252] and generative models [253, 254] which are promising in turbulent flow modeling.
Existing vibrotactile sensation devices such as body suits provide multi-point vibrations but fall short in conveying the sense of fluid flow [33], which is essential for a fully immersive experience. As discussed earlier in Sec. 5, other kinesthetic feedback devices (e.g., exoskeleton) are expensive, less comfortable, and cannot mimic the unique underwater sensations. Therefore, engineering innovations for intuitive haptic sensation is another open problem in this domain.
25 6.4   Latency Mitigation for Long Distance Operation
Mitigating the inherent latency in robot teleoperation has been a focal pursuit in contemporary research, given the pivotal implications these delays impose on overall system performance and operator experience. Researchers have tested USV-ROV pairs with LAN [255] and satellite connection [109] to enable faster communication. For existing tethered ROV systems, video prediction is an exciting research area. While it has gained some attention among the computer vision community [256], it has not been evaluated thoroughly for subsea telerobotics. More specifically, frame interpolation and future frame prediction can be a promising solution for the short-term disruption of video feed.
On the other hand, optical-acoustic dual channels have established wireless communication up to a few hundred meters of depth [257]. RF signals have also been tested for wireless ROV control [258]; however, their application has been limited to shallow water applications. Unlike optical signals, RF waves are undisturbed by water turbulence, turbidity, and solar noise. Therefore, extending the working range of RF signals in a water medium would significantly improve operator-ROV communication in harsh environments. Compact magnetoelectric (ME) antennas [157] are also being explored as a potential modality for robot-robot and robot-teleoperator communications with promising early results.
To account for long-term communication loss, predictive control algorithms and intent-aware robotic systems can offer potential solutions [259]. Predictive control algorithms can utilize historical data and real-time inputs to anticipate future states, while intent-aware systems would enable the robot to switch to autonomous mode and continue performing its tasks when a communication breakdown is imminent. Another approach is to reduce continuous manual input via intelligent delegation of control tasks to the autonomous subsystems within robots and mitigate the impacts of delays on the operator via sensory manipulation [260]. Although this technique has been widely exercised in rehabilitation literature [261], its potential in telerobotics remains under-explored. Further efforts are required from a broader research community that will combine supervisory controls, adaptive algorithms, and advanced sensory feedback into a cohesive framework for addressing the diverse challenges posed by teleoperation delays [260].
26 6.5   Real-World to/from Simulation
Existing virtual simulators are capable of rendering simplified models of underwater workspace, including surface waves [130], chemical profiles [143], coral reefs [142], subsea rigs [134], and some dynamic obstacles [129]. However, they fail to represent the turbulent oceanic conditions and waterbody characteristics [262] in general. As compared earlier in Table 3, only a few simulators (e.g., UWSim and UUV Simulator) provide an extensive choice of sensors and diverse scenarios, making them suitable for testing navigation, SLAM, and path planning algorithms. Many platforms cannot simulate interactions with dynamic marine objects, which is critical for these tasks. The utility comparison of various simulators presented by Ciuccoli et al. [263] also highlights these limitations. Consequently, expanding the sensing capabilities and scenario diversity in simulators remains a pressing area for development.
Another issue is that most existing simulators cannot incorporate real-world data, leaving a gap between simulated scenarios and real-world missions. To address this, comprehensive real2sim frameworks should be developed, capable of translating real-time and historical sensor data from physical deployments into virtual platforms. For instance, project RUMI [28] leverages real-time video feeds to create an interactive terrain within a geo-referenced digital environment. Studies have demonstrated that such frameworks can enhance manipulation and navigation performance through improved realism [264]. Additionally, this approach narrows the gap between virtual and physical models, facilitating sim2real transfer of new algorithms and robotic systems developed in simulated scenarios [265].
Moreover, many existing DT models do not accurately represent their real-world pairs [263]. While they mimic the kinematic properties, onboard sensors and physical properties e.g., buoyancy and center of gravity are not replicated accurately. To improve realism, these design gaps in twin modeling need to be addressed and their interaction with dynamic environments should be extensively compared with real-world cases.
27 6.6   Adaptive Autonomy
The concept of adaptive autonomy goes beyond shared autonomy or supervision. This technology would allow a robot to analyze the operator’s skill level and adjust the degree of autonomy intelligently. Lawrance et al. [35] demonstrate an analogous technology where the robot assesses the operator’s action to provide customized assistance. However, the assessment happens offline, based on the operator’s historical profile; therefore, it cannot provide real-time flexible autonomy during the mission. OceanPlan [48] presents an adaptive holistic re-planner that utilizes sensory feedback to adjust the mission plan in real-time. Researchers have studied adaptive autonomy for vision-based aerial teleoperation [266]. In the future, it can be a breakthrough for long-term subsea operations if the following intelligent and cognitive capabilities are achieved.
First, the robot will analyze which parts of a mission can be handled independently and when operator intervention is necessary. For instance, during long-term navigation, it will engage autopilot mode in calm, open waters, but will prompt the operator to take control in turbulent conditions or when navigating through confined spaces. Second, the robot will be able to detect and correct erroneous commands from the operator. For instance, if an operator-issued command causes an overshoot during object grasping, the robot will override the command and adjust its actions to achieve the ultimate goal. Third, given a high-level objective, the robot will propose multiple solutions, each with attributes such as the shortest route, safest path, or minimum completion time, allowing the operator to select a suitable approach based on mission-specific priorities.
28 6.7   Brain-Machine Interfaces (BMI)
Electrophysiological signals generated from the human brain can be used to control a robot’s motion. To date, most brain-machine/computer interfacing (BMI/BCI) research focuses on assistive technologies for disabled persons, particularly for those who are paralyzed or “locked in" by neurological neuromuscular disorders [267]. The closest parallel work to subsea telerobotics is the remote control of UAVs using BMIs. Nourmohammadi et al. [268] summarize different aspects of SOTA brain-controlled UAV technologies. Other notable BMI applications include construction robots [269], humanoid robots [270], and indoor UGVs [271]. In the underwater domain, Zhang et al. [272] demonstrates a BMI-based simulator, where operators control nine behaviors of a subsea manipulator for grasping marine organisms.
Extensive research is essential to adapt existing BMIs and develop advanced systems for subsea applications. First, a comprehensive analysis of brain waves is necessary to assess their feasibility for guiding complex teleoperation tasks. Second, user cognitive load study is critical to reduce fatigue and ensure that neural control remains effective over long-term missions. Finally, integrating BMI with existing teleoperation interfaces, such as haptic feedback systems and visual displays, needs further study to create a cohesive, multimodal system.
29 7  Conclusions
The growing reliance on ROVs and AUVs for remote subsea expeditions has elevated the importance of designing advanced HMIs to ensure precise control and reliable interaction. Researchers have proposed various interfaces, from early vision-based consoles to more sophisticated, multimodal interfaces that integrate haptics, XR, and NL interactions – to enhance operator efficiency and comfort while equipping robots with greater autonomy and intelligence. In this paper, we reviewed the expansive body of literature on subsea HMI technologies, categorizing and analyzing their strengths, limitations, and applications. We identified unique challenges of subsea teleoperation and SOTA approaches to address them. The role of virtual interfaces and digital twins in lowering development costs, and facilitating operator training is also explored. We further examined SOTA shared autonomy frameworks and their role in bridging operators’ awareness with intelligent robots. Finally, we outlined several promising avenues for future research, including learning from demonstration, human-robot language interaction, cognitively aware systems, improved virtual interfaces, and beyond. Interdisciplinary collaboration across robotics, AI, and human cognitive science is required to meet the unique demands of the underwater domain and further advance HMI systems in subsea telerobotics.
30 Acknowledgments
This work is supported in part by the National Science Foundation (NSF) grants #$23304162330416$ and #$23261592326159$; and the University of Florida (UF) research grant #$132763132763$.

### v2

@@
 Abstract
-This review explores the evolution of human-machine interfaces (HMIs) for subsea telerobotics, tracing back the transition from traditional first-person “soda-straw” consoles (narrow field-of-view camera feed) to advanced interfaces powered by gesture recognition, virtual reality, and natural language models.
-First, we discuss various forms of subsea telerobotics applications, current state-of-the-art (SOTA) interface systems, and the challenges they face in robust underwater sensing, real-time estimation, and low-latency communication.
-Through this analysis, we highlight how advanced HMIs facilitate intuitive interactions between human operators and robots to overcome these challenges.
-A detailed review then categorizes and evaluates the cutting-edge HMI systems based on their offered features from both human perspectives (e.g., enhancing operator control and situational awareness) and machine perspectives (e.g., improving safety, mission accuracy, and task efficiency).
-Moreover, we examine the literature on bidirectional interaction and intelligent collaboration in terms of sensory feedback and intuitive control mechanisms for both physical and virtual interfaces.
-The paper concludes by identifying critical challenges, open research questions, and future directions, emphasizing the need for multidisciplinary collaboration in subsea telerobotics.
+This review explores the evolution of human-machine interfaces (HMIs) in subsea telerobotics, charting the progression from traditional first-person “soda-straw" consoles– characterized by narrow field-of-view camera feeds– to contemporary interfaces leveraging gesture recognition, virtual reality, and natural language processing.
+We systematically analyze the state-of-the-art literature through three interrelated perspectives: operator experience (including immersive feedback, cognitive workload, and ergonomic design), robotic autonomy (contextual understanding and task execution), and the quality of bidirectional communication between human and machine.
+Emphasis is placed on interface features to highlight persistent limitations in current systems, notably in immersive feedback fidelity, intuitive control mechanisms, and the lack of cross-platform standardization.
+Additionally, we assess the role of simulators and digital twins as scalable tools for operator training and system prototyping.
+The review extends beyond classical teleoperation paradigms to examine modern shared autonomy frameworks that facilitate seamless human-robot collaboration.
+By synthesizing insights from robotics, marine engineering, artificial intelligence, and human factors– this work provides a comprehensive overview of the current landscape and emerging trajectories in subsea HMI development.
+Finally, we identify key challenges and open research questions and outline a forward-looking roadmap for advancing intelligent and user-centric HMI technologies in subsea telerobotics.
 keywords:
-Subsea telerobotics; marine robotics; human-machine interface; shared autonomy.
+Human-machine interface; subsea telerobotics.
 1 1. Introduction
-[…]
+Subsea telerobotic technologies have advanced significantly over the past few decades, driven by the need for remote inspection, maintenance, and research expeditions in underwater environments [1, 2].
+Remotely operated vehicles (ROVs) and autonomous underwater vehicles (AUVs) are typically deployed to perform remote tasks in subsea environments that are beyond the reach of human scuba divers [3].
+The robots are generally teleoperated from a surface vessel or a base station [4] for applications such as underwater infrastructure inspection [5, 6, 7], seabed mapping [8, 9], remote surveillance [10, 11], environmental monitoring [12], scientific expeditions, and more.
+Despite growing interest in autonomy, teleoperation continues to dominate subsea missions, comprising over 75% of deployments [13].
+This trend highlights the immense potential of advancing telerobotic interfaces for economic growth and scientific discovery.
 Early teleoperated systems were limited in terms of the feedback they provided, often relying on basic visual data from low-resolution cameras.
@@
 This challenge is especially pronounced in confined or cluttered spaces, a common scenario in underwater pipelines, caves, and shipwreck exploration.
-In larger workspaces, traditional monocular camera-based interfaces fail to provide a comprehensive spatial view, leading to significant cognitive overload on the teleoperators.
-These limitations have led to the development of more interactive systems including 3D scene rendering with augmented/virtual/mixed reality (XR), haptic cues, and natural language (NL) controls for complex teleoperation tasks.
-Recent research highlights the importance of human-robot interface design in improving the operator’s awareness, cognitive load, and stress factors [16, 17, 18] – as well as improving the ROVs’ physical capabilities [19].
-These developments have allowed operators to engage in human-machine shared autonomy, significantly enhancing the precision and safety of subsea missions [20, 21, 22].
-Transparency and reliability are two other critical attributes of shared autonomy, to provide the operator with clear and interpretable insights into the robot’s intentions and actions [23].
-Additionally, fostering mutual trust between the operator and the robot is essential, so that the operator feels both in control and supported by the machine’s autonomous capabilities [24].
-Fig.
-1 presents a general outline of human-machine sensorimotor interaction for subsea telerobotics.
-As shown, the interface elements bridge the human and machine entities with an interactive flow of teleoperation (teleop) control and feedback commands.
-Modern interfaces can integrate virtual engines to render a physics-simulated world and digital twin (DT) models of real ROVs [25, 26].
-These virtual scenarios, simulated independently [27] or modeled from real-world data [28], are useful for mission training and rehearsal, as real underwater deployments involve significant costs and logistics [29].
-DTs also allow for simulating synthetic sensory data and mission logs, supporting the early-stage development and testing of new perception and navigation algorithms in a controlled environment [30, 31, 32].
-Modern subsea ROVs integrate various shared autonomy and semi-autonomous control capabilities, where the robot can handle low-level functions such as station-keeping or obstacle avoidance [33, 22].
-This blend of human awareness with machine intelligence reduces cognitive workload and allows the operator to focus on high-level tasks [34], improving both operational safety and efficiency.
-Customizable interfaces that can adapt to the specific mission needs or the skill level of the operator further streamline the process [35], allowing for greater flexibility and adaptability in a wide range of applications.
+In larger workspaces such as coral reefs, offshore rigs, and mining fields– traditional monocular camera-based interfaces fail to provide a comprehensive spatial view, leading to significant cognitive overload on the teleoperators [16, 17].
 Figure 1: An outline of the human-robot sensorimotor interaction flow for subsea telerobotics is shown.
@@
 Simulators with integrated twin models help teleoperation training by mission rehearsal/replay, which is essential for analyzing teleoperators’ cognitive load as well as for prototyping human-machine embodied/shared autonomy features.
-(FPV: first person view; NL: natural language; XR: augmented, virtual, and mixed reality)
-To this end, contemporary studies emphasize the role of feedback realization, cognitive reasoning, and intuitive control modes for mission safety, efficiency, and success [16, 36].
-Multimodal feedback has been particularly useful in engaging teleoperators’ sensations through means such as vibrotactile cues on the palm [37], force and tension on the forearm [38], augmented visual displays [39], and textual responses to queries [40].
-On the other hand, control commands are typically propagated by traditional buttons, hand gestures [41], body movements [42], or NL instructions [43].
-With any/some of these features, a dynamic feedback-control loop significantly increases the sense of telepresence and enhances mission reliability and success rates of complex subsea tasks.
-The interdisciplinary nature of modern interface technologies, encompassing haptics, XR, and NL features – has led to a diverse body of research literature.
+(FPV: first person view; NL: natural language; XR: augmented/virtual/mixed reality)
+These limitations have led to the development of more interactive systems, including 3D scene rendering with augmented/virtual/mixed reality (XR), haptic cues, and natural language (NL) controls for complex teleoperation tasks.
+Recent research highlights the importance of human-robot interface design in improving the operator’s awareness, cognitive load, and stress factors [18, 19] – as well as improving the ROVs’ physical capabilities [20].
+These developments have allowed operators to engage in human-machine shared autonomy, significantly enhancing the precision and safety of subsea missions [21, 22].
 Fig.
-2 illustrates a snapshot of subsea HMI technologies developed across academia and industry.
-These technologies are often built for specific platforms and applications, revealing an absence of a set of standard guidelines.
-As a result, the body of knowledge remains scattered, with no comprehensive review that compares the technologies, compiles the findings, and highlights opportunities for integration and standardization.
+1 presents a general outline of human-machine sensorimotor interaction for subsea telerobotics.
+As shown, the interface elements bridge the human and machine entities with an interactive flow of teleoperation (teleop) control and feedback commands.
+It increases the sense of telepresence via intuitive feedback, integrates intelligent machine capabilities, and offers virtual engines to render a physics-simulated world and digital twin (DT) model of the real ROV [23].
+More advanced interfaces offer task-sharing between the operator and the robot, thereby reducing cognitive load and enhancing overall mission efficiency [24].
+Knowledge gaps and open questions.
+The interdisciplinary nature of modern HMIs with haptics, XR, and NL features has led to a broad but fragmented body of research literature.
+The existing systems and interfaces are built for a diverse set of platforms and applications without any consistent guidelines.
+There is no comprehensive technical review that compares the interfaces, compiles the findings, and identifies integration opportunities.
+To this end, we pose the following research questions, highlighting key gaps in HMI design, communication modalities, interaction strategies, and autonomy frameworks.
+  1. 1.
+    What design principles and standard guidelines are necessary to ensure modularity and cross-platform compatibility of HMIs in subsea telerobotics?
+  2. 2.
+    What command and feedback modalities (e.g., language, gesture, haptics, camera vision) should be prioritized or fused for different tasks according to complexity, environmental constraints, and mission objectives?
+  3. 3.
+    To what extent can digital replicas (simulators and twins) mimic a real-world scenario, thereby improving operator skills, reducing mission logistics, and supporting algorithm development for subsea operations?
+  4. 4.
+    How various aspects of shared autonomy can be integrated into modern interfaces and adapted according to the operator’s skill level and cognitive load?
+An in-depth understanding of these aspects is essential for driving innovation and advancing the HMI frontiers of subsea telerobotics.
+This paper aims to contribute toward that objective by providing a comprehensive technical review of the key HMI dimensions relevant to marine robotic systems.
+Specific contributions.
+In this paper, we present a structured analysis of subsea HMI systems based on their functional roles, interaction modalities, and levels of autonomy as observed in the literature.
+To facilitate comparative evaluation, we introduce the concept of “technology clusters" that delineate how human-machine interaction happens across different system architectures.
+First, we categorize and evaluate the state-of-the-art (SOTA) subsea telerobotics interfaces and simulators based on a set of guidelines such as their choice of sensory modalities, platform compatibility, user comfort, and optimal use cases (Section 3-4).
+Then, we present a detailed review of human-robot teaming in subsea telerobotics, highlighting various aspects of shared autonomy, bidirectional interaction, and intelligence (Section 5-6).
+Finally, we analyze the SOTA approaches and trends to address the challenges in subsea HMIs, highlighting open problems and use cases for future research (Section 7).
+This review is authored from the perspective of researchers actively involved in subsea robotics to bridge the gap between academic advancements and the practical demands of real-world mission scenarios.
+It aims to provide foundational knowledge to students, researchers, scientists, and engineers engaged in HMI development.
+The classification and open research questions will offer a roadmap to young researchers for systematic investigation into interface design, autonomy integration, and human-robot collaboration.
+For industry leaders, the comparative insights into current systems and emerging trends will inform the design of futuristic HMI solutions.
+2 2. Review Scope And Criteria
+This review spans two primary HMI categories: commercially available interfaces with publicly accessible documentation and peer-reviewed interfaces within the academic literature.
+Although current efforts toward HMI standardization emphasize the adoption of open-source, widely used, and accessible platforms– such as BlueRobotics hardware [25], Robot Operating System (ROS2) middleware [26], and ArduPilot-based autopilot architectures [27]– these initiatives reflect localized, predominantly U.S.-based community practices.
+As such, they do not constitute benchmarks that are consistently adopted across the academic and industrial landscapes.
+Given the lack of established benchmarks in this domain, we introduce a set of evaluation criteria to systematically assess and compare the diverse range of HMI systems.
+Specifically, we formulate a set of guidelines grounded in three core aspects: (i) human operators’ cognitive experience with feedback and control modalities, (ii) robot’s degree of autonomy and their relevance for various tasks, and (iii) the quality of bidirectional human-robot communication.
+Detailed criteria for each evaluation and analysis are presented in the respective sections; itemized features are marked using the following symbols for normalized interpretation:
+  • $\dagger$:  Supported by peer review or official documentation
+  • $\ddagger$:  Derived from demo/promotional video
+[…]
+  • $\blacklozenge$:  Inferred by the authors based on others’ findings; should be treated as illustrative insights rather than definitive claims.
+This approach, while not exhaustive, reflects the broader need for formal standardization efforts (e.g., data formats, open standards, certifications) in HMI technologies.
 Figure 2: A chronological evolution of subsea HMI technologies is shown.
 The top row highlights early achievements and milestones.
-As technology advanced, this century observed industry solutions for interfaces and simulators, incorporating XR and haptic features.
+As technology advanced, this century observed industry solutions for interfaces and simulators, incorporating XR and haptic features (second row).
 In the last five years, NL-based interactions have gained popularity, driven by the success of LLMs.
-In this paper, we compile a comprehensive review of the prominent technologies and relevant literature on subsea HMIs over the last few decades.
-We also identify the knowledge gaps and highlight current trends to facilitate future research.
-Specifically, we make the following contributions.
-  1. 1.
-    We categorize the state-of-the-art (SOTA) subsea telerobotic interfaces and simulators based on attributes such as their choice of sensorimotor interfaces, control and feedback mechanisms, platform compatibility, and other relevant features.
-We then evaluate and compare these features based on their applicability and effectiveness across diverse subsea applications.
-These discussions are in Section 2-3.
-  2. 2.
-    We present a detailed review of human-robot interaction and shared autonomy frameworks in subsea telerobotics, highlighting various design issues regarding bidirectional interaction, degree of intelligence, transparency, trust, and the underlying task allocation processes; see Section 4.
-These analyses also provide valuable insights into human-machine teaming and the associated feedback engineering and interactive control mechanisms; see Section 5.
-  3. 3.
-    Finally, we analyze the SOTA approaches and current trends to address the unique challenges in subsea HMIs, highlighting the open problems and potential new use cases for future research; see Section 6.
-Through this detailed examination of trends, challenges, and potential breakthroughs – we hope to inspire new approaches to bridge the gap between theory and practice.
-We anticipate that this review will serve as a comprehensive resource for researchers and engineers and support the development of more efficient and reliable telerobotic systems to advance subsea robotics technologies of the next generation.
-2 2  Subsea Telerobotic User Interfaces
+3 3. Subsea Telerobotic User Interfaces
 Telerobotic interfaces have evolved significantly over the past decades, with novel integrated features playing a critical role in enhancing operator control and mission success rate.
-Traditionally, vision-based interfaces dominated, where operators relied solely on 2D video feeds to navigate and control a robot [4, 44].
-Haptic feedback was integrated later [37], improving task precision by allowing operators to feel forces such as drag from fluid dynamics and resistance from interactions with objects.
-The introduction of XR interfaces [45] further advanced ROV teleoperation by immersing operators in a 3D virtual environment, providing greater situational awareness and enhancing operator control [46, 47].
-More recently, researchers have integrated natural language understanding capability into mission management interfaces [48], allowing operators to plan missions and control ROVs using NL commands, making interaction more intuitive.
-Table 1: Comparison of existing subsea teleop interfaces based on their integrated features and add-ons. ($\infty$: unlimited number; W: Windows; M: MacOS; L: Linux).
+Traditionally, vision-based interfaces dominated, where operators relied solely on 2D video feeds to navigate and control a robot [4].
+Haptic feedback was integrated later [28], improving task precision by allowing operators to feel forces such as drag and resistance.
+The introduction of XR interfaces further advanced teleoperation by immersing operators in a 3D virtual environment, providing greater situational awareness [29, 30].
+More recently, researchers have integrated natural language understanding capability into interfaces [31], allowing operators to plan missions and control ROVs using NL commands, making interaction more intuitive.
+3.1 3.1   Status Quo of Robot Teleoperation Interface
+Subsea HMIs have been developed to address the unique challenges of underwater operations, where visibility, communication bandwidth, localization techniques, and environmental conditions significantly differ from terrestrial systems.
+Fig.
+2 organizes key milestones in the evolution of HMI technologies into three distinct eras: (i) foundational space and ocean exploration during the 20th century that laid the base for HMI development, (ii) the emergence of marine robotic companies, novel software and hardware solutions in the early 21st century, and (iii) the recent shift toward immersive VR-integrated and language-driven interfaces.
+The motivation behind classifying into these eras is not only due to their chronological relevance but also to highlight how technical, commercial, and human factors have co-evolved in shaping the capabilities and expectations of subsea HMI systems.
+Generally, an ROV teleoperation interface is referred to as ground control station (GCS) [32]; examples of which include QGroundControl [33], APM Planner2 [34], EOD Workspace by GreenSea IQ [35], NaviSuite Mobula [36], etc.
+For smartphones and tablet devices, a few applications are available, including QGroundControl, DroidPlanner [37], SidePilot [38], and MAVPilot [39].
+However, certain interfaces (e.g., SidePilot and MAVPilot) are tailored exclusively for aerial vehicles, relying on GPS signal that is unavailable underwater.
+The diverse construction and application domains of these systems make it essential to categorize and evaluate their functionalities from the operator’s perspective.
+The selected features in Table 1 are chosen to reflect the integration flexibility (e.g., supported OS, open-source availability), built-in safety protocols (e.g., fault diagnostics), and advanced capabilities (e.g., real-time mapping, multi-view 3D displays) for a comprehensive comparison.
+Note that this table is not exhaustive, particularly for commercial interfaces that do not fully disclose their capabilities.
+This table is intended to help readers identify the relevant capabilities and make informed decisions when selecting an interface for their use cases.
+Among open-source interfaces, QGroundControl is the most popular across the ArduPilot community [40].
+It is built on ArduSub [27], a comprehensive open-source framework specifically designed for underwater vehicles.
+Among proprietary interfaces, leading ROV manufacturer VideoRay offers VideoRay Cockpit [41], exclusively designed to operate their ROVs.
+NaviSuite Mobula is a third-party interface built for VideoRay ROVs and offers some autonomous capabilities including scanning a wall, orbiting a target, etc.
+Abyssal Offshore [42] integrates real-time XR visualization, $270\degree270$ ° FOV from multiple cameras, and risk analyses for commercial ROV operation.
+OpenROV cockpit [43] offers browser-based connectivity without requiring any software installation.
+Industry-grade interfaces such as FMC Schilling [44] and Teledyne SeaBotix [45] come as rugged hardware packages and do not require external computers or controllers for teleoperation.
+Table 1: Comparison of existing subsea teleop interfaces based on their feature and platform compatibility.
+($\infty$: unlimited number; W: Windows; M: MacOS; L: Linux); other notations are explained in Sec.
+2.
   Open Video Waypoint Scene 3D Visual Fault Mission Max. # Supported
@@
 DroidPlanner [37] $11$ Android
-3 2.1   Status Quo of Robot Teleoperation Interface
-Subsea HMIs have been developed to address the unique challenges of underwater operations, where visibility, communication bandwidth, localization techniques, and environmental conditions significantly differ from terrestrial systems.
-The interfaces are referred to as ground control stations (GCS) [61]; examples of which include QGroundControl [49], APM Planner2 [51], EOD Workspace by GreenSea IQ [52], NaviSuite Mobula [56], etc.
-For smartphones and tablet devices, a few applications are available including QGroundControl, DroidPlanner [60], SidePilot [62], and MAVPilot [63].
-However, certain interfaces (e.g., SidePilot and MAVPilot) are tailored exclusively for aerial vehicles, relying on GPS signals that are unavailable underwater.
-Table 1 compares some key aspects and available features of SOTA telerobotic interfaces.
-QGroundControl serves as the official interface for submersible (underwater) vehicles within the ArduPilot flight controller developer community [64].
-It is built on ArduSub [65], a comprehensive open-source framework specifically designed for underwater vehicles and integrated into the larger ArduPilot ecosystem.
-Among proprietary interfaces, leading ROV manufacturer VideoRay offers VideoRay Cockpit [53], exclusively designed to operate their ROVs.
-NaviSuite Mobula is a third-party interface built for VideoRay ROVs and offers some autonomous capabilities including scanning a wall, orbiting a target, etc.
-Abyssal Offshore [57] integrates real-time XR visualization, $270\degree270$ ° FOV from multiple cameras, and risk analyses for commercial ROV operation.
-Open-source platforms such as QGroundControl are more suitable for researchers since they allow customization to suit specific hardware configurations and operational needs.
-Moreover, OpenROV cockpit [66] offers browser-based connectivity without requiring any software installation.
-Industry-grade interfaces such as FMC Schilling [67] and Teledyne SeaBotix [54] come as rugged hardware packages and do not require external computers or controllers for teleoperation.
-4 2.2   Vision-based Consoles
-Underwater vision is a well-explored area in telerobotics literature and has long been the key navigational information for human teleoperators [68].
+3.2 3.2   Vision-based Consoles
+Underwater vision is a well-explored area in telerobotics literature and has long been the key navigational information for teleoperators [50].
 The environmental challenges as well as the inherent limitations of 2D visual systems have been studied for decades and many innovative solutions have been proposed.
-We summarize the following challenges and corresponding countermeasures employed in vision-based teleoperation technologies.
-First, traditional consoles with a fixed POV provide limited information to the operator.
-Monocular FPV cameras lack depth perception and fail to provide peripheral vision, resulting in significantly reduced situational awareness.
+We summarize the following challenges and corresponding countermeasures employed in vision-based interfaces.
+First, traditional consoles with a fixed monocular POV lack depth perception and fail to provide peripheral vision, resulting in reduced situational awareness.
 Second, low-light turbid water conditions reduce visibility and hinder the capture of clear images.
-Third, artificial lights used on ROVs to enhance visibility have adverse effects too; their bright light gets reflected and back-scattered by suspended particles directly at the front camera, creating glare and large blind spots for the operator [69].
+Third, ROV’s onboard lights get reflected and back-scattered by suspended particles directly at the front camera, creating glare and large blind spots for the operator [51].
 To compensate for poor visibility and to provide the operator a comprehensive understanding of the scene, modern marine telerobotic systems use high definition wide angle cameras [52], structured lighting systems [53], image enhancement [54], and scene parsing technologies [55].
@@
 Instead, contemporary works present third person view generation process using augmented reality [14], follower ROV [56], external camera [57], etc.
-Although not popular yet, multispectral and hyperspectral imaging has the potential to provide more robust vision systems with enhanced marine life detection and mapping mechanisms [76].
 While vision will remain the fundamental modality for teleoperation, other sensory augmentations are essential for the immersive realization of the remote environment [58].
-Besides underwater ROV’s front camera feed, the visual consoles display a wide range of internal and external sensor readings i.e., telemetry data.
-External sensors capture information such as pressure, depth, temperature, water current, acoustic scans, etc.
-Researchers emphasize scanning sonar systems (e.g.
-side scan sonar, synthetic aperture sonar) for safer navigation, as they detect and display obstacles beyond the limited FOV of cameras [77, 78].
-Internal data encompasses the status of onboard electrical and mechanical systems including energy consumption, housing temperature, data recording, and streaming status, as well as the ROV’s pose and velocity.
-Among these, the attitude and heading reference system (AHRS) estimates and displays the robot’s pose on the console which is crucial for maintaining heading and orientation in feature-deprived underwater scenes [79].
-5 2.3   Integration of XR
+3.3 3.3   Integration of XR
 Figure 3: Snapshots of two telerobotics interfaces are shown; (a) An underwater ROV is being maneuvered with human head movements via HMD and haptic feedback controllers; (b) an AR interface with viewpoint augmentation capability [14] is rendering third-person EOB (eye on the back) views [7] for more peripheral information and situational awareness.
@@
 In $19931993$, NASA launched the first successful Telepresence Remotely Operated Vehicle (TROV) [29] under the Ross Sea ice near the McMurdo Science Station in Antarctica.
-This project demonstrated the feasibility of streaming video feeds from ROV and rendering computer-generated graphics on an HMD.
+This project demonstrated the feasibility of streaming video feeds from ROV and rendering computer-generated graphics on a head-mounted display (HMD).
 Preliminary results also showcased ROV control mechanisms via head movements.
@@
 [59], an XR model should be able to collect, process, transfer, and reconstruct the immersive scene model of the workspace in real-time and enable intuitive robot controls accordingly.
-The cost and logistics associated with long-term underwater missions and the unique environmental complexities pose challenges to achieving these criteria [81].
-Researchers have come up with innovative solutions to address these challenges and have been successful to some extent in XR interfacing between humans and marine robots.
-DOLPHYN [82], a mixed reality-based aquatic interface for multi-sensory exploration of scuba diving sites, allows a diver to remotely operate an ROV while being aware of own diving trajectory.
-Multiple contemporary works explore XR-based view synthesis techniques to provide operators with an enhanced peripheral view during teleoperation.
+The cost and logistics associated with long-term underwater missions and the unique environmental complexities pose challenges to achieving these criteria.
+Researchers have come up with innovative solutions and have been successful to some extent in XR interfacing between humans and marine robots.
+DOLPHYN [60], a mixed reality-based aquatic interface for multi-sensory exploration of scuba diving sites, allows a diver to remotely operate an ROV while being aware of his own diving trajectory.
+Contemporary works explore XR-based view synthesis techniques to provide operators with an enhanced peripheral view.
 For instance, the augmented visualization of the CoMAS project [61] supports ROV pilots during underwater archaeological site exploration and artifacts collection.
@@
 Windowed or stereoscopic VR [29] further improves perception, when compared against monoscopic VR or desktop-based visuals [17].
-Other VR-based frameworks such as TWINBOT [86, 87] offer full teleoperation experiences in simulation, which is particularly useful for pilot training prior to real-world missions.
+Other VR-based frameworks such as TWINBOT [64] offer virtual teleoperation, which is particularly useful for pilot training.
 The impact of training and XR visuals is supported by studies showing a reduction in task completion time by over 50%, even when the control method remains unchanged [17].
-Two key components of XR systems: simulators and digital twins – are further discussed in Sec.
-3.
-6 2.4   Multimodalities Interface
-Besides vision, human sensorimotor control relies on multimodal sensory feedback such as visual, auditory, and somatosensory (tactile and proprioceptive) cues to interpret the consequences of initiated actions [88, 89, 90].
-Haptic feedback is the most recognized one in enhancing XR systems; researchers have explored innovative methods for simulating tactile sensations and haptic devices are employed to deliver haptotactile stimulation (e.g., vibrations and force feedback) synchronized with real-time events [91, 92].
-Haptic-integrated interfaces typically feature a visual feed via a desktop or a head mounted display (HMD) and one/multiple haptic feedback devices such as wearable gloves, exoskeletons, body suits, etc.
+Simulators and digital twins for XR systems are further discussed in Sec.
+4.
+3.4 3.4   Multimodalities Interface
+Besides vision, human sensorimotor control relies on multimodal sensory feedback such as auditory and somatosensory (tactile and proprioceptive) cues to interpret the consequences of initiated actions [65].
+Haptic feedback is the most recognized one in enhancing XR systems; researchers have explored innovative methods for simulating haptotactile stimulation (e.g., vibrations and force feedback) synchronized with real-time events [66].
+Haptic-integrated interfaces typically feature a visual feed via a desktop or an HMD and one/multiple haptic feedback devices such as wearable gloves, exoskeletons, body suits, etc.
 Preferred choices for haptic feedback include vibrotactile, skin indent, pressure, etc.
 on finger and wrist regions as these options only require cheap hardware compared to exoskeletons or full-body haptic suits.
-Fig.
-3 shows a pair of hand-held controllers for vibrotactile feedback.
-Other haptic feedback mechanisms and their suitable applications are categorized in Table 2.
-The benefits of haptic feedback have been extensively studied for ROV driving and manipulation tasks [93, 94].
+To clusterize SOTA haptic feedback mechanisms, we define the criteria and terminology used in Table 2 as follows.
+  • Feedback location and type represent the anatomical region and sensory modality used to provide physical stimulus to the operator.
+While multiple locations and feedback types can be integrated in a single system, the classification is based on the dominant feedback point and function.
+  • Control mechanism captures the physical console and sensor suites responsible for interpreting and delivering the operator’s command/intention to the machine.
+  • Ideal use cases are set by reviewing the user studies and demonstrations reported in literature; metrics such as task relevance and accuracy improvement are considered for this purpose.
+Note that the examples are not exclusive, rather illustrative of commonly reported subsea applications.
+  • Ergonomics reflects a qualitative estimation of the interface’s ease of use and fatigue factors.
+This is judged based on user studies (where available) and author-reviewed qualitative descriptions from cited works.
+The benefits of haptic feedback have been extensively studied for ROV driving and manipulation tasks [67].
 Ryden et al.
-[2] present a framework for assisted manipulation (opening and closing of a valve) using real-time haptic feedback to the operator’s hand.
-It guides the robotic arm to stay away from forbidden regions using force feedback and pulls the arm toward the desired orientation.
-Such force feedback is crucial for the operator’s sense of virtual presence and is driven by different sources such as artificial muscle [95], magneto-rheological fluid [96, 97], and dc motors [38].
-Researchers also explore how other external sensory information can be incorporated into the feedback system.
-For instance, Amemiya et al.
+[2] present a framework for assisted valve manipulation using real-time haptic feedback to the operator’s hand.
+It guides the robotic arm to stay away from forbidden regions and pulls toward the desired orientation using force feedback.
+To incorporate other external sensory feedback, Amemiya et al.
 [68] develop a system that combines water pressure and torsion forces to enhance kinesthetic perception.
-Moreover, the haptic interface developed by Xia et al.
-[13, 99] includes the feedback of hydrodynamic state as a vector field in addition to other vibrotactile cues, helping the operator to maintain orientation and heading of the ROV in turbulent water.
-Furthermore, whole body motion mapping frameworks [42, 100] take advantage of body dexterity and movements to control multiple motions of an ROV.
+The haptic interface developed by Xia et al.
+[58, 69] includes the feedback of hydrodynamic state as a vector field, helping the operator to maintain orientation and heading of the ROV in turbulent water.
+Whole body motion mapping frameworks [70, 71] take advantage of body dexterity and movements to control multiple motions of an ROV.
 However, as mentioned in Table 2, haptic feedback on multiple sensory organs may reduce user comfort [72].
@@
 Additionally, guiding systems are developed for haptic-supported manipulation that evaluate discomfort in human hand using the Rapid Upper Limb Assessment (RULA) [73], and direct the hand toward a more comfortable posture [74].
-Overall, despite the limitations of current haptic devices and the low ergonomics in some applications, the combination of visual and haptic feedback have been demonstrated significant potential to enhance the intuitive sense of telepresence.
-Table 2: Haptics and XR capabilities of telerobotic interfaces are categorized and their use cases are highlighted.
+Overall, despite current limitations in haptic devices and ergonomics, visuo-haptic feedback has demonstrated significant potential to enhance the intuitive sense of telepresence.
+Table 2: Haptics and XR capabilities of telerobotic interfaces are categorized and their use cases are highlighted. Criteria are defined in Sec. 3.4, notations are explained in Sec. 2.
   Feedback Feedback Control Ideal Use Cases User Comfort Selected
@@
   Fingertips Vibrotactile Wearable gloves, Grasping, collecting sample, High [28, 2]
-stylus controlling valve/knob [104, 105]
+stylus controlling valve/knob [75]
 Wrist, Force Wearable gloves, Sensing drag and tension, Medium [30, 76, 77]
@@
 force, fluid flow actuated chair using dexterity for manipulation
-7 2.5  Integration of Natural Language
+3.5 3.5  Integration of Natural Language
 NL-based interfaces leverage the natural human way of communicating and reasoning for more intuitive mission planning and control.
@@
 Classic visual language systems such as RoboChatTag [84] and RoboChatGest [85] use sequences of symbolic patterns to convey simple instructions from diver to robot/operator, utilizing AR-Tag markers and hand gestures, respectively.
-Recently, natural language processing (NLP) with deep learning networks has allowed the smooth conversion of spoken/written language into understandable and actionable commands for robotic systems [116].
-Early NL interfaces focus on basic navigation commands or status queries such as “take a left", “report battery status" [117], and have since evolved to support more sophisticated mission planning and contextual understanding.
+Recently, natural language processing (NLP) with deep learning networks has allowed the smooth conversion of spoken/written language into actionable commands for robotic systems [86].
+Early NL interfaces focus on basic navigation commands or status queries such as “take a left", “report battery status" [87], and have since evolved to support more advanced contextual understanding.
 Fig.
-4 provides a chronological overview of NL integration in telerobotic control and mission management.
-Figure 4: A timeline progression of natural language integration in subsea telerobotic interfaces and mission management tasks over the past $1515$ years is shown.
+4 provides a chronological overview of NL functionalities in telerobotic control and mission management.
+Figure 4: Progression of natural language functionalities in subsea telerobotic interfaces and mission management tasks over the past $1515$ years is shown.
 The key works referenced are: Hallin et al.
@@
 These interfaces operate as offline planners, processing full mission descriptions and planning the optimum trajectory before deployment [88].
+Figure 5: An example of a language-guided mission planning interface is shown [94]; verbal instructions are parsed into sub-tasks such as S (start), Mv (move), Sp (spiral), Az (adjust), E (end) – to generate deployable mission waypoints.
 The planner first converts spoken or written instructions into structured commands using speech recognition and syntactic parsing.
-This is followed by semantic analysis, where the system interprets the meaning and maps it to specific robotic actions or a sequence of tasks.
+This is followed by semantic analysis, where the system interprets the meaning and maps it to specific robotic actions.
 A particular example for Word2Wave interface [94] is shown in Fig.
@@
 Such interfaces enable a robotic system to plan and execute complex missions from high-level operator instructions.
-In contrast to offline planning interfaces, online interfaces allow real-time dialogue between the operator and
-Figure 5: An example of a language-guided mission planning interface [43].
-The robot parses verbal instructions into sub-tasks such as S (start), Mv (move), Sp (spiral), Az (azimuth), E (end); and then generates mission waypoints.
-the robot.
+In contrast to offline planning interfaces, online interfaces allow real-time dialogue between the operator and the robot.
 Hastie et al.
-[119] demonstrate an online dialogue module named REGIME (Report Generation from Metadata), where sensor data and event logs are summarized into mission updates upon operator demand.
+[89] demonstrate REGIME dialogue where sensor data and event logs are summarized into mission updates upon operator demand.
 Robb et al.
 [90] introduce a chat-based interface for controlling a group of robots, allowing the operator to inquire about the status of individual robots via the leader and adjust their behavior as needed.
-This reduces the cognitive load on operators by shifting focus from managing multiple robots to a single control point through the leader.
+This reduces the operator’s cognitive load by shifting focus from managing multiple robots to a single control point through the leader.
 In ChatSim [92], dialogue capabilities are utilized to customize a simulated environment.
 The operator first asks for specific structures (e.g.
-canyon, shipwreck, coral reef, oyster) to be added into the virtual world, followed by issuing commands to the robot such as “go through the canyon and find the shipwreck”, “move in a circle of radius $33$ m over the oysters”, etc.
-The current challenges with integrating language models into telerobotic systems are multifaceted.
-First, the computational demands of LLMs may exceed the capabilities of resource-constrained underwater systems.
-To address this, researchers are developing smaller language models (SLMs) while technologies such as LoRA [123] are helping to adapt LLMs for real-time applications in edge devices.
-Second, ensuring that the robot accurately interprets commands and understands spatial references requires sophisticated models, extensive training data, and robust error-handling mechanisms [116].
-The consequences of misinterpretation can be severe in high-risk subsea operations, hence current research focuses on making these systems more resilient, context-aware, and capable of handling complex, multi-step instructions [124].
-Moreover, online commands can be error-prone due to variations in dialect, tone, or pronunciation [125].
-Instead, an offline planning interface allows the operator to review and correct for potential errors before launch.
+canyon, shipwreck, coral reef, oyster) to be added into the virtual world, followed by issuing navigation commands such as “go through the canyon and find the shipwreck”.
+However, online commands can be error-prone due to variations in dialect, tone, or pronunciation [95].
+Instead, an offline planning interface allows the operator to review and correct potential errors before launch.
 Further discussions on this are presented in Sec.
-6.
-Figure 6: A digital twin of an underwater ROV is shown in action; (a) a human is teleoperating it to pick up an object using head movements and finger grips; (b) the twin model rendered in HoloOcean is mimicking the action in virtual environment.
-8 3 Simulators and Digital Twins
-Simulation environments offer a valuable alternative for prototyping, error debugging, and pilot training for marine robotic systems in a controlled, risk-free setting, without requiring expensive logistics and access to real waterbodies.
-These simulators not only replicate the dynamics of underwater conditions but also support the integration of digital replicas of physical systems, facilitating the testing and development of robust control algorithms and mission planning strategies for real robots.
-Fig.
-6 shows an example of underwater grasping with a digital twin rendered in the HoloOcean simulator.
-9 3.1  Digital Twins (DTs)
-A digital twin refers to a simulated replica of a real robot that possesses similar attributes and can mimic the motion of its real-world pair within a virtual environment [126, 127].
+7.1.
+Figure 6: A digital shadow of an underwater ROV is shown in action; (a) a human is teleoperating it to pick up an object using head movements and finger grips; (b) the replica rendered in HoloOcean is mimicking the action in virtual environment.
+4 4. Simulators and Digital Twins
+Simulators offer a valuable alternative for prototyping, error debugging, and pilot training for marine robotic systems in a controlled, risk-free setting, without requiring expensive logistics and access to real waterbodies.
+They not only replicate the environmental dynamics but also support the integration of digital replicas of physical systems, facilitating the development of control algorithms and mission planning strategies.
+4.1 4.1  Digital Twins (DTs)
+A digital twin refers to a simulated replica of a real robot that possesses similar attributes and can mimic the motion of its real-world pair within a virtual environment [96].
 The concept of DT for field robots was first presented by NASA, where the digital entity was defined as an integrated multi-physics, multi-scale, probabilistic simulation of a vehicle or system that mirrors the life of its real-world peer [97].
-This project envisioned four major applications of the DT technology.
-  • Pre-mission simulation: The twin will simulate the vehicle’s future missions before launch, enabling studies of mission parameters, anomalies, and mitigation strategies to enhance mission success.
-  • Real-time flight mirroring: The twin will continuously update and mirror the actual flight with real-time sensor data, predicting flight status for ongoing analysis and adjustment.
-  • In-situ forensics: The twin will allow diagnosis and analysis of potential catastrophic faults in real-time and address anomalies during the flight.
-  • Mission modification analysis: The twin will assess the impact of unforeseen mission parameter changes (e.g., failed actuators), and analyze new load distributions to inform mission decisions.
-DTs exhibit varying levels of complexity, ranging from basic motion replication to comprehensive emulation of physical and operational attributes.
-The simplified setup shown in Fig.
-6 does not replicate the real testbed in the simulator; however, the digital robot mimics the motion of the real ROV, thereby qualifying as its twin.
-For more sophisticated DTs, the digital replica must mirror key physical properties (e.g.
-dimensions, weight, buoyancy), energy consumption profile, sensor configurations, and interaction with the external environment.
-Examples of some SOTA simulators and their DT collections include BlueROV2 in UNav-Sim [129], Triton XLX in PaleBlue [130] and VMAX simulator [131], VideoRay Pro5 and Defender in Era’s VideoRay simulator [132], and Girona500 in UWSim [133]; see Fig.
-7.
+This project envisioned four applications of the DT technology including pre-mission simulation, real-time modification, monitoring, and onboard forensics for comprehensive analyses.
+Although two terms – digital shadow and twin – are used interchangeably in literature, there is a clear distinction between them.
+A shadow is a passive digital replica that mirrors the real-time entity without active predictive modeling or feedback loops [98].
+It typically supports one-way (i.e., real2sim) data flow, and is limited to basic monitoring.
+A true DT would include high-fidelity representations of vehicle dynamics, sensor fusion models, environmental interactions, and a bi-directional data pipeline enabling predictive simulation [99].
+For instance, the simplified setup shown in Fig.
+6 does not offer predictive simulation; however, the digital robot mimics the motion of the real ROV, thereby qualifying as its shadow rather than a twin.
 Figure 7: Snapshots of a few marine robotic simulators are shown; the top row shows proprietary simulators for professional use [100, 101, 102]; the bottom row shows open source simulators mainly used for academic research [23, 103, 104].
 The use of DTs has been reported in different branches of robotics including aerial, manufacturing, surgical, space, and marine robots.
-All DT systems are built upon some common backends with domain-specific techniques for data acquisition and rendering [127].
 Generally, the raw sensor data acquired from the robot is processed with physics engines and reconstructed as a high-fidelity DT model [96].
@@
 Instead, during a communication disruption, the predictive nature of a DT system can be utilized to create a realistic representation as if communication had remained uninterrupted.
-DT systems have also been deployed in underwater sensor networks to test communication reliability and synchronization accuracy between multiple nodes [136].
 More recently, Adetunji et al.
 [108] integrate SLAM and path planning capabilities in a versatile DT system.
-Other applications of DT in marine robotics include pipeline inspection [137], assistive control [13], trajectory planning [30], etc.
-10 3.2  Underwater Simulators
+Other applications of marine DT include inspection [109], assistive control [58], trajectory planning [106], etc.
+4.2 4.2  Underwater Simulators
 A simulator is essential to work with DTs as well as to mimic the external environment parameters.
@@
 Independently designed virtual ROVs or DTs of real ROVs are rendered in these platforms to test their kinematic properties and interactions with the surroundings.
-Researchers have studied different underwater vehicles such as gliders [30, 139, 140], ROVs [135, 26], AUVs [137], underwater legged robots [141], and humanoid robots [111] in these simulators.
+Researchers have studied different underwater vehicles such as gliders [106, 111], ROVs [107, 23], AUVs [109], and humanoid robots [81] in these simulators.
 Yang et al.
@@
 Hu et al.
-[139] model a blended-wing-body underwater glider (BWBUG) and investigate the accuracy of a pitch controller under the influence of ocean currents.
+[111] model a blended-wing-body underwater glider and investigate the accuracy of a pitch controller under ocean current influence.
+Table 3 compares SOTA underwater simulators developed across academia and industry.
+Among proprietary simulators, Abyssal [101] and ACSM [112] offer a rich collection of underwater infrastructures and work-class ROV models, designed to train operators in mining, repairing, and maintenance of offshore rigs.
+PaleBlue [100] and VMAX Simulator [102] allow users to import custom-designed ROVs from CAD models and simulate their behavior, forces, frictions, etc.
+in the virtual world.
+Era Marine and EIVA have partnered with VideoRay to create advanced simulators for the VideoRay Pro5 and Defender ROVs.
+These simulators provide advanced tools for planning, surveying, and underwater inspections.
+M8-R Sim [113], originally developed for the MATE ROV Competition, has over $140140$ simulation scenarios, including those for monitoring the growth of Crown of Thorns Starfish (COTS).
+Unlike higher-end professional simulators, open-source alternatives offer fewer built-in scenarios and ROV models but allow for better customization and development according to mission-specific needs.
+For instance, UWSim [104], StoneFish [114], and HoloOcean [103] support ROS with sensor plugins e.g., side scan sonar, imaging sonar, acoustic beacon, DVL, magnetometer, force sensor, etc.
+UUV Simulator [23] integrates chemical particles and a simulation scenario to measure their concentration using a CPC (condensation particle counter) sensor.
+To summarize, open-source solutions are better suited for research and development with ROS support, whereas licensed products offer superior training opportunities for commercial ROV operation.
 Table 3: Comparison of underwater simulators are shown.
-Basic state estimation sensor plugins (e.g., IMU, GPS, depth) are common for all simulators, therefore omitted here; indicates digital twin of a real ROV.
+Basic state estimation sensor plugins (e.g., IMU, GPS, camera, depth) are common for all simulators, therefore omitted here; : digital twin of a real ROV, other notations are explained in Sec.
+2.
 (a) Comparison based on sensor integration capability, scenario libraries, and available robotic agents.
-  Sensor Plugin Built-in Scenarios Robot Models
-  PaleBlue [130] RGB-D camera, laser Reefs, oil and gas rigs Triton XLX
-Abyssal [134] Camera Rigs, sand cloud, pipelines Abyssal ROV
-M8-R Sim [142] Camera 140+ coral reefs with starfishes M8-R ROV
-Era VideoRay [132] Camera, sonar, lights 6 seabed scenarios with shipwrecks VideoRay Pro5 , Defender
-GRi VROV [67] Camera and sonar Pipelines and rigs LittleGeek, BigGeek
-UUV Simulator [26] Camera, sonar, DVL, CPC Lakes, shipwrecks, coast, BOP panels A9 , LAUV , Saga , RexROV
-UUVSim [143] Camera, sonar Open water, current, seabed 6T UUV, RexROV
-HoloOcean [27] Camera, sonar, beacon, DVL Dams, pier harbors, pipelines, open water 6 ROV agents
-UWSim [133] Camera, DVL, multi-beam sonar Dredging, pipelines, seabed with amphora Girona500
-UNav-Sim [129] Camera, LiDAR Seabed with algae and sand-covered pipelines BlueROV2 heavy
-MARUS [144] Camera, LiDAR, sonar, DVL, AIS Ships, human divers, seabed One AUV
-StoneFish [145] RGB-D camera, sonar, DVL Sea surface, seabed, wind, sun Girona500
-NaviSuite [56] Camera, sonar, DVL, USBL Cables, pipelines, and ship anchors VideoRay Defender
-VMAX Simulator [131] Camera, sonar Dynamic seabed, cables, ship with crane Triton , Comanche , Mojave
-ACSM [146] Camera Rigs and pipelines Triton XLX
-Unicom [147] FPV and TPV cameras Pipelines and rigs with valves, knobs, doors 2 work class ROVs
+  Sensor Plugin Built-in Scenarios Robot Models Optimal Usage
+  PaleBlue [100] Laser Reefs, oil and gas rigs Triton XLX  VR-supported rig inspection
+Abyssal [101] - Rigs, sand cloud, pipelines Abyssal ROV Manipulator control
+M8-R Sim [113] - 140+ coral reefs with starfishes M8-R ROV Education & novice training
+Era VideoRay [115] Sonar, lights 6 scenarios with shipwrecks VideoRay Pro5 , Defender  Training for VideoRay ROVs
+GRi VROV [44] Sonar Pipelines and rigs LittleGeek, BigGeek Maintenance, rescue, mine detection
+UUV Simulator [23] Sonar, DVL, CPC Shipwrecks, coast, BOP panels A9 , LAUV , Saga , RexROV Perception & control study
+UUVSim [116] Sonar Open water, current, seabed 6T UUV, RexROV Hydrodynamic study, DL training
+HoloOcean [103] Sonar, beacon, DVL Dams, pier harbors, pipelines 6 ROV agents Sim2Real, multi-agent control
+UWSim [104] DVL, multi-beam sonar Dredging, pipelines, amphora Girona500  Planning, control, HIL study
+UNav-Sim [117] LiDAR Algae, sandy pipelines BlueROV2 heavy  Navigation & inspection study
+MARUS [118] LiDAR, sonar, DVL, AIS Ships, human divers, seabed One AUV Acoustic sensing study
+StoneFish [114] Sonar, DVL Sea surface, seabed, wind, sun Girona500  Localization & navigation study
+NaviSuite [36] Sonar, DVL, USBL Cables, pipelines, and ship anchors VideoRay Defender  Inspection, intervention
+VMAX Simulator [102] Sonar Cables, ship with crane Triton , Comanche , Mojave  Manipulator control
+ACSM [112] - Rigs and pipelines Triton XLX  Deployment & manipulator control
+Unicom [119] - Rigs with valves, knobs, doors 2 work class ROVs Manuevering risk assessment
 (b) Comparison based on the availability of various features and supported platforms.
@@
 OS (W,M,L) ✓✕✕ ✓✕✕ ✓✕✓ ✓✓✓ ✓✕✕ ✓✓✓ ✓✕✓ ✓✓✓ ✓✕✓ ✓✕✓ ✓✓✓ ✓✕✕ ✓✕✕ ✓✕✕ ✓✓✓
-Fig.
-7 illustrates some SOTA underwater simulators developed across academia and industry.
-Among proprietary simulators, Abyssal [134] and ACSM [146] offer a rich collection of underwater infrastructures and work-class ROV models, designed to train operators in tasks such as underwater mining, pipeline repairing, and maintenance of offshore oil and gas rigs.
-PaleBlue [130] and VMAX Simulator [131] allow users to import custom-designed ROVs from CAD models and simulate ROV behavior, forces, frictions, etc.
-in the virtual world.
-Era Marine and EIVA have partnered with VideoRay to create advanced simulators for the VideoRay Pro5 and Defender ROVs.
-These simulators, Era’s VideoRay [132] and EIVA NaviSuite [56], provide advanced tools for planning, surveying, and underwater inspections.
-M8-R Sim [142], originally developed for the MATE ROV Competition, has over $140140$ simulation scenarios, including those for monitoring and estimating the growth of Crown of Thorns Starfish (COTS).
-Unlike higher-end professional simulators, open-source alternatives offer fewer built-in scenarios and ROV models but allow for better customization and further development according to mission-specific needs.
-For instance, UWSim [133], StoneFish [145], and HoloOcean [27] support numerous sensors e.g., side scan sonar, single beam and multi-beam imaging sonar, acoustic beacon, DVL, magnetometer, force sensor, etc.
-UUV Simulator [26] integrates chemical particles and a simulation scenario to measure their concentration using a CPC (condensation particle counter) sensor.
-They also provide closed water (lake) and open water (coast) scenarios, each with different current profiles.
-As shown in Table 3, the open-source solutions are better suited for research and development with ROS support, whereas licensed products offer superior training opportunities for commercial ROV operation.
-11 4 Human-Machine Shared Autonomy
-Telerobotic systems are categorized into three types according to their intelligent decision-making capabilities: (i) direct teleoperator control with no intelligence, (ii) shared/collaborative reasoning, and (iii) autonomous/self-sufficient systems.
+5 5. Human-Machine Shared Autonomy
+Subsea robotic systems are categorized into three types according to their intelligent decision-making capabilities: (i) direct teleoperator control with no intelligence, (ii) shared/collaborative reasoning, and (iii) autonomous/self-sufficient systems.
 In direct teleoperation, the operator has total control over the robot’s motion and actions, generally through a live feed and a motion controller.
@@
 These robots rely on advanced algorithms for perception, decision-making, and execution, making them ideal for repetitive tasks in familiar environments.
-However, their utility is constrained by environmental uncertainties, the cost of logistics, and the limitations of current AI-based technologies [149, 150].
-12 4.1  Direct Teleoperation Mode
+However, their utility is constrained by environmental uncertainties, the cost of logistics, and the limitations of current AI-based technologies [122].
+5.1 5.1  Direct Teleoperation Mode
 Direct or hands-on teleoperation involves controlling the movements of an underwater robot using low-level commands without intelligent decision-making by the robot.
@@
 It allows for responsive control, enabling precise adjustments that automated systems may struggle with.
-Despite high cognitive demand, direct teleoperation remains the preferred approach for manipulation tasks involving high risk and fine precision, where the operator controls one joint angle at a time in a joint-by-joint fashion using a traditional console [151]; examples include wreck inspection [152], valve control [153], benthic sampling [151], etc.
+Despite high cognitive demand, direct teleoperation remains the preferred approach for manipulation tasks involving high risk and fine precision, where the operator controls one joint angle at a time in a joint-by-joint fashion [123]; examples include structure inspection [124], valve control [125], benthic sampling [123], etc.
 These systems are challenged by the inherent communication delays in underwater environments, which can reduce control precision and responsiveness.
-The delays further complicate tasks that depend on real-time feedback [154, 153].
-To address this, researchers have developed adaptive control strategies to compensate for both model uncertainties and external disturbances during teleoperation [155, 153].
-Furthermore, electromagnetic and optical communication modes offer promising solutions for establishing high-speed reliable links [156, 157], improving the real-time commands and feedback flow in direct teleoperation systems.
+To address this, researchers have developed adaptive control strategies to compensate for both model uncertainties and external disturbances during teleoperation [125].
+Furthermore, electromagnetic and optical communication modes offer promising solutions for establishing high-speed reliable links [126], improving the real-time commands and feedback flow in direct teleoperation systems.
 Table 4: The table summarizes various types of intelligent machines and their autonomous task-sharing capabilities reported across the subsea telerobotics literature.
+Criteria are defined in Sec.
+5.2, notations are explained in Sec.
+2.
   Type of Human (Teleoperator) Side Machine (Underwater ROV) Side Selected
@@
 Multi-robot Operate the Medium Coordinate and control Coverage, inspection, [131, 9]
-collaboration leader ROV follower ROVs search and rescue [6, 162]
+collaboration leader ROV follower ROVs search and rescue [6]
 Mission Set mission Low Generate waypoints, Exploration, mapping, [91, 31]
@@
 Learning from Demonstrate Initially high Learn knowledge base, Routine maintenance, [132, 133]
-demonstration relevant tasks adapt to novel tasks sample collection [104, 165]
-Context-aware Provide contextual Medium Search for relevant features, Inspection, surveillance, [5, 166]
-exploration cues or guidance report key findings environmental monitoring [167]
-Docking Guide the robot Low Identify docking point, Charging, data transfer, [168, 169]
-assistance close to the dock self-correct fine motion repair, recovery [170, 171]
-13 4.2   Human-Robot Collaborative Reasoning
-The goal of collaborative decision-making systems is to offer a blend of human oversight along with automated functions.
-These systems allow operators to set high-level tasks or objectives while the robot handles low-level operations, thus reducing cognitive load on the operator and enabling more efficient operation [22].
-For instance, an operator can command the simple autonomous task of hovering; the robot determines its state using an IMU, DVL, and depth sensors to correct for any drift motion to remain stationary.
+demonstration relevant tasks adapt to novel tasks sample collection [134]
+Context-aware Provide contextual Medium reason a context, Inspection, surveillance, [5, 124]
+exploration cues or guidance plan accordingly environmental monitoring
+5.2 5.2   Human-Robot Collaborative Reasoning
+The goal of collaborative decision-making systems is to offer a blend of human oversight along with automated functions [135].
+These systems allow operators to set high-level objectives while the robot handles low-level operations, thus reducing operator’s cognitive load and enabling more efficient operation [22].
+For instance, an operator can issue hovering command; the robot determines its state using an IMU, DVL, and depth sensors to correct for any drift motion to remain stationary.
 More advanced reasonings are achieved in a supervisory manner (for a single robot) or with a leader-follower strategy (for a group of robots).
-Table 4 provides an overview of the SOTA intelligent capabilities of underwater robots and their task-sharing mechanisms with humans.
-Mission planning, online mission adjustment, and fault diagnosis systems benefit from shared autonomy while keeping the operator in the loop [172], but without demanding continuous oversight.
-Research in supervised planning and navigation demonstrates that a new mission path suggested by a circumstance-aware robot is often safer and reduces task completion time [34, 20, 48, 173].
+Table 4 aims to present a structured overview of current capabilities and highlight how autonomy is distributed in SOTA subsea HMI systems.
+The clustering is motivated by a need to understand and compare the evolving roles of humans and machines in underwater missions.
+In modern telerobotic systems, autonomy is no longer binary (teleop vs.
+autonomous) but exists on a spectrum.
+Thus, the table categorizes task-sharing models by considering: (i) the nature of human intervention, (ii) the degree of machine intelligence, and (iii) the operational context (e.g., planning, fault handling, coordination).
+The cognitive load classification is informed by findings from user studies and qualitative insights reported by other researchers.
+For instance, mission planning, online mission adjustment, and fault diagnosis systems benefit from shared autonomy while keeping the operator in the loop [136], but without demanding continuous oversight.
+Research in supervised planning and navigation demonstrates that a new mission path suggested by a circumstance-aware robot is often safer and reduces task completion time [128, 31, 137].
 The GALILEO planning interface adjusts the course of actions and resource allocations based on high-level goals set by the operator [138].
-Additionally, intelligent robots are capable of detecting and reporting internal faults [175] and data anomalies [159], reducing the need for constant operator oversight.
-Such distributed control also allows the operator to focus on mission objectives while offloading low-level tasks to the robot.
+Additionally, intelligent robots are capable of detecting and reporting internal faults and data anomalies [129], reducing the need for constant operator oversight.
 Sharing scientific objectives with a context-aware robot is beneficial too, as the robot can search for specific cues or features under operator guidance.
@@
 SAUVIM (1998199819981998[140] is one of the earliest semi-autonomous vehicles with a 6-DOF manipulator arm developed for deep water intervention tasks.
-More recently, supervisory control frameworks such as DexROV [109] and SHARC [21] have been developed where the operator performs remote manipulation tasks using language commands and hand gestures while experiencing the task in immersive virtual reality [34].
+More recently, supervisory control frameworks such as DexROV [79] and SHARC [24] have been developed where the operator performs remote manipulation tasks using language commands and hand gestures [128].
 The GRASPER project [141] utilizes a pre-scanned scene point cloud to inform the robot about obstacle locations.
@@
 Lee et al.
-[20] propose a shared teleoperation method where humans can provide real-time input by sketching a desired path on the visual console, guiding the intelligent system, which plans an optimum path along that drawn trajectory and executes the motion.
+[21] propose a shared teleoperation method where humans sketch a desired path on the visual console, guiding the intelligent system, which plans an optimum path along that drawn trajectory and executes the motion.
 The adaptive supervision system built within the GreenSea interface provides operators customized assistance based on their skill levels [120].
@@
 Their approach achieves a ten-fold reduction in manipulation task completion time using a simulated DexROV; however, they did not consider vehicle dynamics or external disturbances encountered in real-world environments.
-Among real-world projects, PANDORA [164] demonstrates the role of LfD in automating inspection and intervention tasks at subsea oil and gas facilities.
+Among real-world projects, PANDORA [133] demonstrates the role of LfD in automating intervention tasks at subsea oil and gas facilities.
 The authors emphasize the necessity of a cognitive knowledge base enriched with semantic information about the workspace, as opposed to relying solely on pre-programmed trajectories or task-specific descriptions.
@@
 These followers possess some local autonomy and operate as a coordinated group.
-The human operator, controlling the leader, oversees the collective movement of the fleet [183].
-Two fundamental challenges in such fleets are formation control and trajectory control [184].
+The human operator, controlling the leader, oversees the collective movement of the fleet.
+Two fundamental challenges in such fleets are formation and trajectory control [146].
 The operator plays a crucial role in modulating the fleet’s velocity through the leader to ensure the followers maintain formation, avoid collisions, and stay synchronized with the leader.
-Such robotic fleets, also known as swarms, are useful for tasks requiring extended area coverage, dense scans, and rapid completion.
-Notable applications include search and rescue [161], structure inspection [6, 185], ship hull inspection [186, 187], large area mapping [9, 188, 189], cooperative localization [190], and continuous surveillance of underwater facilities [162].
+Such robotic fleets, also known as swarms, are useful for tasks requiring extended area coverage and rapid completion.
+Notable applications include search and rescue [131], structure inspection [6], ship hull inspection [147], large area mapping [9], and cooperative localization [148].
 The complexity of these tasks exceeds the capabilities of a single vehicle or imposes an excessive mental load on the operator.
@@
 [149] characterizes operators as independent from the swarm, limiting their role to mission initialization and maintenance; however, this perspective does not cover the full spectrum of multi-robot literature.
-For instance, several studies [192, 193] explore scenarios where follower robots autonomously estimate their relative pose while the operator directly controls the leader.
+Several studies [150, 151] present scenarios where follower robots autonomously estimate their relative pose while the operator directly controls the leader.
 Moreover, contemporary research emphasizes heterogeneous robot groups, unlike traditional homogeneous swarms, where all robots had the same characteristics and role [152].
-In heterogeneous settings, the operator allocates tasks based on each robot’s specific capabilities and real-time operational status [195].
-For marine applications, heterogeneous groups typically include an unmanned surface vehicle (USV) that acts as a communication relay between ROVs and the operator; an example is demonstrated in [6], consisting of one ASV and two SWIM-R ROVs for shallow water pipeline inspection.
-These configurations enhance mission reliability by leveraging the USV as an additional wireless communication channel.
-14 4.3   Autonomous Systems
-As opposed to direct teleoperation and shared autonomy schemes discussed above, autonomous systems or AUVs operate independently, requiring no control commands from the operator.
+In heterogeneous settings, the operator allocates tasks based on each robot’s specific capabilities and operational status [153].
+Heterogeneous groups typically include an unmanned surface vehicle (USV) that acts as a communication relay between ROVs and the operator; an example is demonstrated in [6], consisting of one ASV and two SWIM-R ROVs for shallow water pipeline inspection.
+These configurations enhance mission reliability by leveraging the USV as an additional communication channel.
+5.3 5.3   Autonomous Systems
+As opposed to direct teleoperation and shared autonomy schemes, autonomous systems or AUVs operate independently, requiring no control commands from the operator.
 Subsequently, HMI for an AUV is primarily designed for status updates and mission management rather than direct control.
-However, even fully autonomous systems rely on human operators for mission planning, progress monitoring, and emergency intervention.
-This highlights the critical role of AUV interface design in ensuring the safety and effectiveness of autonomous missions.
+However, even fully autonomous systems rely on human operators for mission planning, monitoring, and intervention.
 Generally, the interface includes a software console, a surface control module (SCM), and a transducer to maintain real-time communication with the AUV.
@@
 Kaeli et al.
-[159] develop an anomaly detection framework that enables AUVs to focus on high-resolution scanning and transmit only the anomalous data to the operator for inspection, reducing data congestion and improving operational efficiency.
-A similar approach is reported for crack detection in subsea pipelines using a camera, laser, and multi-beam echosounder (MBES); the AUV detects potential cracks and sends the scanned images to an operator for further review [5].
+[129] develop an anomaly detection framework that enables AUVs to focus on high-resolution scanning and transmit only the anomalous data to the operator, reducing data congestion and improving efficiency.
+A similar approach is reported for crack detection in subsea pipelines using a camera, laser, and multi-beam echosounder; the AUV detects potential cracks and sends the scanned images to an operator for further review [5].
 The operator then requests close-up detailed examination if necessary, or continues the mission.
-Commercial AUV/UUV Platforms.
+Current status & limitations.
 Currently, the top application segments for AUV/UUVs are military and defense, offshore oil and gas drilling, oceanography, archaeology, and alternative energy harvesting [155].
@@
 Consequently, access to AUVs has been restricted to governments and their agencies, industry leaders, and a handful of academic and research labs.
-Examples of SOTA AUVs are L3Harris IVER (US), A9-E ECA (France), Teledyne Gavia (US), Kongsberg Hugin (Germany), SAAB Seaeye Sabertooth (Sweden), Exail A18 (France), and Atlas Elektronik SeaCat (Germany).
-Another lightweight low-cost category known as micro-AUV is gaining popularity for academic and research purposes; examples include NemoSens (France), Seaber Yuco series (France), and EcoSUB (UK).
-A detailed review of current AUV technologies, their challenges, and applications is outlined in [197].
-15 5  Engineering Integration of Interactive Control
+Notable platforms such as Kongsberg Hugin (Germany), SAAB Seaeye Sabertooth (Sweden), and Exail A18 (France) have demonstrated long-range deployments for pre-scripted missions with limited surface intervention.
+Military-grade AUVs such as L3Harris IVER (US) and Atlas Elektronik SeaCat (Germany) offer surveying, intelligence gathering, mine disposal, etc.
+A lightweight, low-cost category known as micro-AUV is gaining popularity for academic and research purposes; examples include NemoSens (France), Seaber Yuco series (France), and EcoSUB (UK).
+However, true autonomy – involving decision-making, adaptive planning, and high-level semantic understanding – remains an ongoing research challenge.
+We summarize the following limitations by reviewing the SOTA AUV literature.
+  • Environmental uncertainty: Underwater environments are highly dynamic, with varying turbidity, current, and visibility, making perception and long-term localization unreliable.
+  • Communication and energy constraints: Lack of real-time power and data transfer prevents operator supervision and long-term deployments.
+  • Limited contextual reasoning: Current autonomous systems cannot reason about high-level goals or modify behavior based on task complexity or environmental cues.
+  • Fault tolerance: Unlike teleoperated systems, AUVs lack immediate human oversight, which raises concerns for risk assessment and fault recovery.
+Development trends.
+Recent academic and industrial trends in AUVs reflect a push toward scalability, adaptability, and improved intelligence.
+One notable direction is multi-AUV (swarm) systems for cooperative task execution (e.g., distributed mapping, search and rescue), which significantly enhance efficiency and spatial coverage [156].
+Demand for low-cost and low-power AUVs is on the rise for academic research [157].
+Additionally, learning-based techniques (e.g., LfD) are substituting explicit programming, as they allow AUVs to acquire complex skills by observing human operators [158].
+Finally, the fusion of DT models with AUV platforms is gaining traction for risk mitigation, allowing for virtual planning and testing of mission strategies.
+6 6. Engineering Integration of Interactive Control
 Figure 8: Examples of bidirectional human-robot interaction in telerobotic tasks are shown, highlighting modern control and feedback mechanisms.
@@
 8 demonstrates some examples of human-robot interaction through command and feedback exchanges during teleoperated missions.
-16 5.1  Visual Feedback Technique
+6.1 6.1  Visual Feedback Technique
 Visual feedback is typically sent from cameras mounted inside a pressure-sealed housing, providing a front-facing first-person view of the robot.
@@
 Augmented third-person views, when combined with real-time data overlays such as sonar or lidar maps [162, 172], further render a rich virtual environment for interaction.
-17 5.2   Haptic Feedback Technique
+6.2 6.2   Haptic Feedback Technique
 Haptic feedback involves sensory information derived from mechanoreceptors embedded in the skin (tactile input) and muscles, tendons, and joints (kinesthetic inputs), provided to the human operator [173].
@@
 Furthermore, the force-generating actuators tend to be large and power-intensive, leading to tactile feedback being the preferred choice for most applications.
-18 5.3   Pseudo-Haptic Feedback
+6.3 6.3   Pseudo-Haptic Feedback
 Given the high costs and logistical challenges of open water ROV deployment, virtual environments are beneficial for prototyping and user training.
@@
 [58], and SHARC-VR [24].
-19 5.4   Other Feedback: Logs, Maps, Auditory Cues
+6.4 6.4   Other Feedback: Logs, Maps, Auditory Cues
 In addition to visual and haptic feedback, teleoperators receive information about the robot’s surroundings through other sensors.
@@
 Nevertheless, the authors point out that auditory feedback has much lower spatial resolution compared to vision, as humans are far more adept at visual navigation [196].
-20 5.5   Interactive Control
+6.5 6.5   Interactive Control
 The existing works explore various mechanisms to issue control commands to ROVs, each with unique strengths depending on the task and the operational environment.
@@
 […]
-21 6 Challenges and Open Problems
+7 7. Challenges and Open Problems
 The subsea telerobotics domain presents unique challenges due to the hostile and dynamic nature of the underwater environment.
 Key technical hurdles include limited FOV, latency in long-distance operations, hydrodynamic variations, uncertainty in state estimation, and the presence of dynamic obstacles within confined spaces.
-Open problems in this sector span the fields of robotics, low-light perception, AI, computer vision, human cognition, and ocean engineering.
-Fig.
-12 highlights the promising directions for future research.
-Few prominent research areas and their open problems are discussed below.
+Open problems in this sector span the fields of robotics, low-light perception, AI, computer vision, human cognition, and ocean engineering; see Fig.
+12.
 Figure 12: Key open research problems in the subsea telerobotics are summarized.
 Operator-centric research emphasizes enhancing telepresence through immersive feedback and effortless control strategies, while ROV-focused research prioritizes advanced sensing capabilities and intelligent decision-making.
-22 6.1   Human-Machine Dialogue in Natural Language
+7.1 7.1   Human-Machine Dialogue in Natural Language
 As natural language has been proven to be one of the most effortless control methods for teleoperation, the future goal is to enable real-time natural conversations, where the robot would possess a cognitive understanding of the goal and actively engage in decision-making.
-LLMs, particularly GPT models [237, 238], have successfully conducted real-time human-machine dialogues for various HRI settings [239, 240].
-However, marine robots face the additional challenges of integrating sensory data and mission objectives into their decision-making process while executing a teleoperator’s command.
+LLMs, particularly GPT models [204], have successfully conducted real-time human-machine dialogues for various HRI settings [205, 206].
+However, marine robots face the additional challenges of integrating sensory data into their decision-making process while executing an operator’s command.
 ChatSim [92] and OceanGPT [93] have shown some progress in this regard, demonstrating limited dialogue capabilities and gathering vast knowledge of the ocean environment into the model, respectively.
-Further research can enable the robot to understand various instructions, reason logically, and propose alternative strategies using a cognitive understanding of a mission.
+Further research can enable robots to reason logically and propose alternative strategies.
 While language models can parse instructions and reason from a long conversation, they struggle with linguistic biases [95] such as variations in dialect/pronunciation, indirect commands, as well as mission-specific terminologies.
-Therefore, another research direction would be a risk assessment for voice-operated subsea missions and developing fail-safe mechanisms for the robot to identify, query, and correct ambiguous or potentially dangerous commands.
-23 6.2   Optimizing Learning from Demonstration (LfD)
+Therefore, another research direction would be a risk assessment for NL interfaces and developing error-handling mechanisms for the robot to identify, query, and correct ambiguous or potentially dangerous commands.
+In natural language generation (NLG) domain, errors are corrected via human feedback and reinforcement learning [207].
+Lin et al.
+[208] compare three error handling strategies– reprompt, suggestion, and confirmation– across users of different age groups interacting with a chatbot.
+Among telerobotic NL interfaces, ChatSim [92] utilize closed perception-action loops to minimize error in GPT-generated mission plan.
+However, other interfaces are yet to adopt any structured error-handling mechanisms.
+To clarify ambiguous user queries, researchers use Similar Question Model (SQM) [209] that identifies ambiguous utterances by extracting features from past similar queries.
+Other approaches to reduce dialect and linguistic error include phonetic normalization, code mixing (i.e., mixing words from multiple language) [210] etc.
+To this end, these correction methods have been applied to search agents and AI assistants [211], hence their utility for robotic interfaces remains underexplored.
+7.2 7.2   Optimizing Learning from Demonstration (LfD)
 Robot’s ability to learn from repetitive tasks and apply that knowledge to unknown scenarios remains under-explored in marine robotics, unlike terrestrial and aerial domains [212].
@@
 However, extensive real-world evaluation is necessary to adopt these developments into industrial applications.
-In general, any reinforcement learning (RL) pipelines struggle with dynamic changes such as robot’s component failure or unseen terrain [242].
+In general, reinforcement learning (RL) pipelines struggle with dynamic changes such as robot’s component failure or unseen terrain [213].
 Extensive training (i.e., demonstrations) in diverse scenarios is a straightforward solution.
@@
 Meta RL [214] and online RL [213] have shown progress in such model adaptation through reward-driven optimization.
-Researchers have utilized these approaches for improving subsea locomotion [244], path planning [245], and target tracking [246].
-Second, realistic simulators featuring various underwater scenes can be a potential alternative for recording numerous task demonstrations.
-Third, algorithmic studies are needed to determine the minimum number of demonstrations required to achieve a desired performance threshold.
-Additionally, hybrid approaches that combine LfD and RL should be explored, where operator interventions during autonomous missions are treated as demonstrations or corrective feedback.
-Such interventions could potentially teach the robot to generalize knowledge, refine policies, and avoid similar errors in the future.
-24 6.3   Hydrodynamic Sensing & Haptic Feedback Modeling
+Researchers have utilized these approaches for improving subsea locomotion [215] and path planning [216].
+Second, realistic simulators featuring various scenes can be a potential alternative for recording numerous demonstrations.
+Third, algorithmic studies can reveal the minimum number of demonstrations required to achieve a desired performance threshold.
+Additionally, hybrid approaches (LfD and RL) should be explored, where operator interventions during autonomous missions are treated as demonstrations or corrective feedback.
+Such interventions could potentially teach the robot to generalize knowledge and refine policies.
+7.3 7.3   Hydrodynamic Sensing & Haptic Feedback Modeling
 Modeling haptic sensations from remote underwater settings presents multi-faceted challenges.
 First, the extreme pressure and complex hydrostatic/hydrodynamic forces limit the sensing accuracy [58].
-Second, existing fluid dynamic models cannot accurately represent detailed information for different spatiotemporal scales of turbulent water flows.
-Third, translating high-fidelity hydrodynamic data at both the micro and macro levels in intuitive haptic feedback that effectively engages teleoperators’ sensory perception while maintaining ergonomics – remains as open problem.
-In recent years, researchers have explored materials such as Polycrystalline Nickel Titanium [247], Carbon Nanotubes [248], and Tungsten Disulfide Nanosheets [249], which exhibit flexible, hydrophobic, and corrosion-resistant properties.
+Second, existing fluid dynamic models cannot accurately represent different spatiotemporal scales of turbulent water flows.
+Third, translating high-fidelity hydrodynamic data at both micro and macro levels in intuitive haptic feedback that engages operators’ sensory perception while maintaining ergonomics – remains an open problem.
+In recent years, researchers have explored new sensing materials such as Polycrystalline Nickel Titanium [217] and Tungsten Disulfide Nanosheets [218], which exhibit hydrophobic and corrosion-resistant properties.
 Doppler current profiler and artificial lateral line sensors have shown promising results for far-field and near-field flow measurements, respectively [69].
 Other techniques such as Smoothed Particle Hydrodynamics (SPH) [219] and Position Based Dynamics (PBD) [220] are being explored for modeling free surface flows.
-Potential research directions in this regard are physics-driven learning [252] and generative models [253, 254] which are promising in turbulent flow modeling.
-Existing vibrotactile sensation devices such as body suits provide multi-point vibrations but fall short in conveying the sense of fluid flow [33], which is essential for a fully immersive experience.
-As discussed earlier in Sec.
-5, other kinesthetic feedback devices (e.g., exoskeleton) are expensive, less comfortable, and cannot mimic the unique underwater sensations.
-Therefore, engineering innovations for intuitive haptic sensation is another open problem in this domain.
-25 6.4   Latency Mitigation for Long Distance Operation
-Mitigating the inherent latency in robot teleoperation has been a focal pursuit in contemporary research, given the pivotal implications these delays impose on overall system performance and operator experience.
-Researchers have tested USV-ROV pairs with LAN [255] and satellite connection [109] to enable faster communication.
-For existing tethered ROV systems, video prediction is an exciting research area.
-While it has gained some attention among the computer vision community [256], it has not been evaluated thoroughly for subsea telerobotics.
-More specifically, frame interpolation and future frame prediction can be a promising solution for the short-term disruption of video feed.
-On the other hand, optical-acoustic dual channels have established wireless communication up to a few hundred meters of depth [257].
+Potential research directions in this regard are physics-driven learning [221] and generative models [222], which show promise in turbulent flow modeling.
+7.4 7.4   Latency Mitigation for Long Distance Operation
+Mitigating the inherent latency in robot teleoperation has been a focal pursuit in contemporary research, given its impact on overall system performance and operator experience.
+Researchers have tested USV-ROV pairs with LAN [223] and satellite connection [79] for faster communication.
+For existing tethered ROV systems, video prediction and frame interpolation is an exciting research areas.
+While it has gained some attention among the computer vision community [224], it has not been evaluated thoroughly for telerobotics.
+On the other hand, optical-acoustic dual channels have demonstrated wireless communication up to a few hundred meters of depth [225].
 RF signals have also been tested for wireless ROV control [226]; however, their application has been limited to shallow water applications.
@@
 To account for long-term communication loss, predictive control algorithms and intent-aware robotic systems can offer potential solutions [227].
-Predictive control algorithms can utilize historical data and real-time inputs to anticipate future states, while intent-aware systems would enable the robot to switch to autonomous mode and continue performing its tasks when a communication breakdown is imminent.
-Another approach is to reduce continuous manual input via intelligent delegation of control tasks to the autonomous subsystems within robots and mitigate the impacts of delays on the operator via sensory manipulation [260].
+Predictive control utilizes historical data and real-time inputs to anticipate future states, while intent-aware systems enable the robot to switch to autonomous mode and continue performing its tasks when a communication breakdown is imminent.
+Another approach is to reduce continuous manual input via intelligent delegation of control tasks to mitigate the impacts of delays on the operator via sensory manipulation [228].
 Although this technique has been widely exercised in rehabilitation literature [229], its potential in telerobotics remains under-explored.
-Further efforts are required from a broader research community that will combine supervisory controls, adaptive algorithms, and advanced sensory feedback into a cohesive framework for addressing the diverse challenges posed by teleoperation delays [260].
-26 6.5   Real-World to/from Simulation
+Further efforts are required to combine supervisory controls, adaptive algorithms, and advanced sensory feedback into a cohesive framework for addressing the challenges posed by teleoperation delays [228].
+7.5 7.5   Real-World to/from Simulation
 Existing virtual simulators are capable of rendering simplified models of underwater workspace, including surface waves [100], chemical profiles [116], coral reefs [113], subsea rigs [101], and some dynamic obstacles [117].
 However, they fail to represent the turbulent oceanic conditions and waterbody characteristics [230] in general.
-As compared earlier in Table 3, only a few simulators (e.g., UWSim and UUV Simulator) provide an extensive choice of sensors and diverse scenarios, making them suitable for testing navigation, SLAM, and path planning algorithms.
-Many platforms cannot simulate interactions with dynamic marine objects, which is critical for these tasks.
-The utility comparison of various simulators presented by Ciuccoli et al.
-[263] also highlights these limitations.
-Consequently, expanding the sensing capabilities and scenario diversity in simulators remains a pressing area for development.
-Another issue is that most existing simulators cannot incorporate real-world data, leaving a gap between simulated scenarios and real-world missions.
-To address this, comprehensive real2sim frameworks should be developed, capable of translating real-time and historical sensor data from physical deployments into virtual platforms.
-For instance, project RUMI [28] leverages real-time video feeds to create an interactive terrain within a geo-referenced digital environment.
-Studies have demonstrated that such frameworks can enhance manipulation and navigation performance through improved realism [264].
-Additionally, this approach narrows the gap between virtual and physical models, facilitating sim2real transfer of new algorithms and robotic systems developed in simulated scenarios [265].
-Moreover, many existing DT models do not accurately represent their real-world pairs [263].
-While they mimic the kinematic properties, onboard sensors and physical properties e.g., buoyancy and center of gravity are not replicated accurately.
-To improve realism, these design gaps in twin modeling need to be addressed and their interaction with dynamic environments should be extensively compared with real-world cases.
-27 6.6   Adaptive Autonomy
+As compared earlier in Table 3, only a few simulators (e.g., UWSim and UUV Simulator) provide an extensive choice of sensors and diverse scenarios to support navigation, SLAM, and planning algorithms.
+Many platforms cannot simulate interactions with dynamic marine objects, which is critical for the aforementioned tasks.
+The utility comparison of various simulators [231] also highlights these limitations.
+Expanding the sensor suites and scenario diversity in simulators remains an open research area.
+Another limitation of existing simulators is their inability to incorporate real-world data, leaving a gap between simulation and real-world missions.
+Robust real2sim frameworks are needed to translate real-time and historical sensor data into virtual environments.
+For instance, project RUMI [166] leverages real-time video feeds to create an interactive terrain within a digital environment to improve realism [232].
+Eventually, this approach may facilitate sim2real transfer of algorithms developed in simulation [233].
+Moreover, many digital shadow/twin models fail to accurately replicate real-world pairs, lacking fidelity in sensor modeling and environmental dynamics [231].
+Addressing these gaps is essential for achieving operational realism and system performance of simulated environments.
+7.6 7.6   Adaptive Autonomy
 The concept of adaptive autonomy goes beyond shared autonomy or supervision.
@@
 Third, given a high-level objective, the robot will propose multiple solutions, each with attributes such as the shortest route, safest path, or minimum completion time, allowing the operator to select a suitable approach based on mission-specific priorities.
-28 6.7   Brain-Machine Interfaces (BMI)
+7.7 7.7   Brain-Machine Interfaces (BMI)
 Electrophysiological signals generated from the human brain can be used to control a robot’s motion.
@@
 Finally, integrating BMI with existing teleoperation interfaces, such as haptic feedback systems and visual displays, needs further study to create a cohesive, multimodal system.
-29 7  Conclusions
-The growing reliance on ROVs and AUVs for remote subsea expeditions has elevated the importance of designing advanced HMIs to ensure precise control and reliable interaction.
-Researchers have proposed various interfaces, from early vision-based consoles to more sophisticated, multimodal interfaces that integrate haptics, XR, and NL interactions – to enhance operator efficiency and comfort while equipping robots with greater autonomy and intelligence.
-In this paper, we reviewed the expansive body of literature on subsea HMI technologies, categorizing and analyzing their strengths, limitations, and applications.
-We identified unique challenges of subsea teleoperation and SOTA approaches to address them.
-The role of virtual interfaces and digital twins in lowering development costs, and facilitating operator training is also explored.
-We further examined SOTA shared autonomy frameworks and their role in bridging operators’ awareness with intelligent robots.
-Finally, we outlined several promising avenues for future research, including learning from demonstration, human-robot language interaction, cognitively aware systems, improved virtual interfaces, and beyond.
-Interdisciplinary collaboration across robotics, AI, and human cognitive science is required to meet the unique demands of the underwater domain and further advance HMI systems in subsea telerobotics.
+8 8. Conclusions
+The increasing dependence on ROVs and AUVs for subsea operations underscores the critical need for advanced HMIs that enable precise control, robust interaction, and mission success.
+In this review, we synthesized a broad spectrum of literature on subsea HMI technologies, systematically categorizing and analyzing their capabilities, limitations, and application domains.
+We identified the unique operational challenges inherent to subsea teleoperation and examined state-of-the-art approaches designed to mitigate them.
+Special attention was given to the role of virtual interfaces and digital twins in reducing development costs and enhancing operator training.
+Furthermore, we explored shared autonomy frameworks that integrate operator intent with autonomous decision-making to improve situational awareness and system responsiveness.
+The review concludes by highlighting several promising directions for future research, including learning from demonstration, natural language-based interaction, and cognitively adaptive systems.
+Addressing these challenges will require sustained interdisciplinary collaboration across robotics, artificial intelligence, and cognitive science to advance the next generation of HMI systems for subsea telerobotics.
 9 Acknowledgments